### v1

# Structural Properties of Entropic Vectors and Stability of the Ingleton Inequality

Abstract
We study constrained versions of the Ingleton inequality in the entropic setting and quantify its stability under small violations of conditional independence. Although the classical Ingleton inequality fails for general entropy profiles, it is known to hold under certain exact independence constraints. We focus on the regime where selected conditional mutual information terms are small (but not zero), and the inequality continues to hold up to controlled error terms. A central technical tool is a structural lemma that “materializes” part of the mutual information between two random variables, implicitly capturing the effect of infinitely many non-Shannon–type inequalities. This leads to conceptually transparent proofs without explicitly invoking such infinite families. Some of our bounds recover, in a unified way, what can also be deduced from the infinite families of inequalities of Matúš (2007) and of Dougherty–Freiling–Zeger (2011), while others appear to be new.
1 Introduction
In this paper we study inequalities for Shannon entropy, and, in particular, variants of the Ingleton inequality in the entropic setting. It is well known that the Ingleton inequality in its original form does not hold for entropies of certain tuples of jointly distributed random variables; see, e.g. [Mat95, Csi96, ZY98, Mat99, Ham+00]. However, under certain restrictions (or with suitable additional terms) the inequality may still hold. We discuss several manifestations of this important phenomenon.
We begin with a brief review of the standard framework for information inequalities. For a tuple of random variables $(X_{i}\;{\bm{:}}\;i=0,\dots,n-1)$ its entropy profile is the function on the power set of the set $\left\{0,\dots,n-1\right\}$ indexing the tuple. The value of this function on $J\in 2^{[n]}$ is the entropy of the joint distribution $(X_{i}\;{\bm{:}}\;i\in J)$. Since the entropy of the empty collection is always zero, we drop the corresponding value from consideration. Thus, every entropy profile is effectively a point in $\mathbb{R}^{2^{n}-1}$. In particular, for a four-tuple of random variables, its entropy profile is a 15-dimensional vector.
A point in $\mathbb{R}^{2^{n}-1}$ is called entropic if it is the entropy profile of the probability distribution of some tuple $(X_{i}\;{\bm{:}}\;i=0,\dots,n-1)$. The set of all entropy profiles for all $n$-tuples of random variable is usually denoted $\Gamma_{n}^{\star}$. The points in the closure of the set of entropy profiles are called almost entropic points. This set is denoted by $\bar{\Gamma}_{n}^{\star}$. It has been shown in [ZY97] that this set is in fact a convex cone. The geometry of this cone is quite non-trivial. It is known, for example, that for all $n\geq 4$ the set $\bar{\Gamma}_{n}^{\star}$ is not polyhedral, [Mat07]). An exact description of this cone for $n\geq 4$ is still unknown. However, there are some known lower and upper bounds for this set. In what follows we briefly discuss the basic facts about $\bar{\Gamma}_{4}^{\star}$.
An entropy profile of a four-tuple of jointly distributed random variables can be thought of as a function assigning a real number to every non-empty subset of $\{0,1,2,3\}$. It is known that every such function $\varphi$ necessarily satisfies the inequalities expressing the properties of non-negativity, monotonicity, and submodularity:
$0\leq\varphi(A)\leq\varphi(B)$ $\text{for all sets}\ A\subset B\subset\left\{0,1,2,3\right\},$
$\varphi(A\cup B)+\varphi(A\cap B)\leq\varphi(A)+\varphi(B)$ $\text{for all sets}\ A,B\subset\left\{0,1,2,3\right\}.$
In other words, every entropy profile is a non-negative, monotone, submodular function. In information theory, these constraints for $\varphi$ are usually called Shannon inequalities. In algebra, the functions that respect these constraints are usually called rank functions of a polymatroid. This is a fairly standard and well-studied class of objects.
The set of all $\varphi$ satisfying these inequalities is traditionally denoted $\Gamma_{4}$ (the cone of submodular functions on a four-point set). This set is clearly a superset (i.e., an upper bound) of $\bar{\Gamma}_{4}^{\star}$.
Some sufficient conditions for being an entropic point can be formulated with the help of the Ingleton inequalities. In what follows we use the notation: for a set function $\varphi$ define
$\varphi(x:y)$ $:=\varphi(\left\{x\right\})+\varphi(\left\{y\right\})-\varphi(\left\{x,y\right\}),$
$\varphi(x:y|z)$ $:=\varphi(\left\{x,z\right\})+\varphi(\left\{y,z\right\})-\varphi(\left\{x,y,z\right\})-\varphi(\left\{z\right\}).$
With this notation, the Ingleton inequality can be written in the following compact form:
$\varphi(0:1)\leq\varphi(0:1|2)+\varphi(0:1|3)+\varphi(2:3).$ (1)
Permuting the set $\left\{0,1,2,3\right\}$ we obtain $\binom{4}{2}=6$ similar inequalities that differ in the order of the variables. We denote by $\Gamma^{\rm Ing}_{4}$ the cone of all non-negative, monotone, submodular functions $\varphi$ on $\left\{0,1,2,3\right\}$ that satisfy, in addition to the Shannon inequalities, the six versions of the Ingleton inequality on a four-point set. This set is a subset (lower bound) of the set of almost entropic points: it is known that every $\varphi$ in $\Gamma^{\rm Ing}_{4}$ is almost entropic, see [MS95] (independently, this was proved in [Ham+00]).
The sets $\Gamma_{4}$ and $\Gamma_{4}^{\mathrm{Ing}}$ are both polyhedral cones, and the set of entropic points is sandwiched between the two: $\Gamma_{4}^{\mathrm{Ing}}\subsetneq\bar{\Gamma}_{4}^{\star}\subsetneq\Gamma_{4}.$ The closure of $\Gamma_{4}\setminus\Gamma_{4}^{\mathrm{Ing}}$ is the union of six closed convex cones with disjoint interiors. They are symmetric under permutations of the set $\left\{0,1,2,3\right\}$. From now on we consider just one of these cones in which the strict version of inequality (1) is violated, and we refer to it as the non-Ingleton cone.
The non-Ingleton cone has a 14-dimensional simplex as its base. One of the 13-dimensional faces of this simplex is defined by the Ingleton equality and is a common face with the base of $\Gamma_{4}^{\mathrm{Ing}}$. The remaining vertex of the simplex is the so-called special point. It is neither an entropic nor an almost entropic point, as was shown in [ZY98]. The study of the cone of almost entropic points essentially reduces to the question of what neighborhood around the special point is free of almost entropic points, or, equivalently, what neighborhood of the face shared between the Ingleton and non-Ingleton regions is filled with almost entropic points. [^1] If we denote by $\alpha_{1},\dots,\alpha_{15}$ the convex coordinates in the non-Ingleton simplex (the base simplex of non-Ingleton cone), such that $\alpha_{15}=1$ is the defining equation of the special point, while $\alpha_{15}=0$ is the equation of the face common with the Ingleton-region, then the almost-entropic region is described by the inequality $\alpha_{15}\leq\Phi(\alpha_{1},\dots,\alpha_{14})$ where $\Phi$ is some concave function on the $13$-dimensional simplex. Ultimately, the complete characterization of $\bar{\Gamma}^{*}_{4}$ reduces to the description of the function $\Phi$. The main results of this paper contribute to a better understanding of the behavior of $\Phi$ near the boundary of its domain. A detailed discussion of the geometry of the cones $\bar{\Gamma}_{4}^{\star}$, $\Gamma_{4}^{\mathrm{Ing}}$, and $\Gamma_{4}$, and of the non-Ingleton cone can be found, e.g., in [Csi16, Csi25] and in [MC16, MP19]. In general, the complete description of $\bar{\Gamma}_{4}^{\star}$ remains far from being understood.
Constraint versions of the Ingleton inequality.
In what follows we discuss the Ingleton inequality for a four-tuple of random variables. To avoid ambiguity and improve the readability of the notation we denote the four-tuple of jointly distributed random variables $(X,Y,A,B)$, and rewrite (1) with the usual notation for the conditional mutual information as
$I(X:Y)\leq I(X:Y|A)+I(X:Y|B)+I(A:B)$ (ING)
We will consider several cases in which some of the quantities
I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X),I(X:Y|A),I(X:Y|B)I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X),I(X:Y|A),I(X:Y|B) (2)
are small compared with other information quantities. It is known that (ING) holds for joint distributions $(X,Y,A,B)$ satisfying certain independence constraints, that is, whenever some of the quantities in (2) vanish. More specifically, the following implications are known:
[…]
as well as the symmetric statements obtained by exchanging $A$ with $B$. These conditional constraints were stated explicitly in [KR13], although they were implicitly known earlier, since they follow directly from the infinite families of inequalities established by Matúš in his seminal paper [Mat07].
We focus on the possible violation of the Ingleton inequality in the regime where some of the mutual information terms in (2) are small but still positive. We consider three cases:
  1. (i)
    $\max\{I(X:A|Y),\,I(A:Y|X)\}\leq\varepsilon\,I(X:Y)$;
  2. (ii)
    $\max\{I(X:A|Y),\,I(X:Y|A)\}\leq\varepsilon\,I(A:Y)$
  3. (iii)
    $\max\{I(A:X|Y),\,I(A:Y|X),\,I(B:X|Y),\,I(B:Y|X)\}\leq\varepsilon\,I(X:Y)$.
If only two conditional mutual informations are bounded, as in cases (i) and (ii), then the Ingleton inequality holds for such a four-tuple up to an additive error of order $\sqrt{\varepsilon}$. These cases are addressed in Theorem A and Theorem B. In contrast, when all four conditional mutual informations are bounded, as in case (iii), the Ingleton inequality remains valid up to an additive error of order $\varepsilon\log\varepsilon^{-1}$, as in Theorem C below.
Theorem A (name=,restate=thsqrt).
Let $(X,Y,A,B)$ and $\varepsilon\geq 0$ be such that
$\max\left\{I(A:X|Y),I(A:Y|X)\right\}\leq\varepsilon\cdot I(X:Y).$
Then
$I(X:Y)\leq I(X:Y|A)+I(X:Y|B)+I(A:B)+O(\sqrt{\varepsilon})\cdot I(X:Y).$
Theorem B (name=,restate=thsqrtii).
Let $(X,Y,A,B)$ and $\varepsilon\geq 0$ be such that
$\max\left\{I(X:Y|A),I(A:Y|X)\right\}\leq\varepsilon\cdot I(X:A).$
Then
$I(X:Y)\leq I(X:Y|A)+I(X:Y|B)+I(A:B)+O(\sqrt{\varepsilon})\cdot I(X:A).$
Theorem C (name=,restate=theloge).
Let $(X,Y,A,B)$ be a four-tuple of random variables such that
$\max\{I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X)\}\leq\varepsilon\cdot I(X:Y).$
Then
$I(X:Y)\leq I(X:Y|A)+I(X:Y|B)+I(A:B)+O(\varepsilon\cdot\log\varepsilon^{-1})\cdot I(X:Y).$
Proofs of Theorems A and B are given in Section 4, Theorem C is proven in Section 6.
Remark 1.
A special case of Theorem A, the implication (3), is the essentially conditional inequality (${\cal I}4^{\prime}$) discussed in [KR13]. From [KR13, theorem 3 (claim 4)] it follows that the term $O(\sqrt{\varepsilon})$ in this theorem cannot be replaced by a function of order smaller than $O(\varepsilon^{2/3})$. Closing the remaining gap between the upper bound $O(\sqrt{\varepsilon})$ and the lower bound $\Omega(\varepsilon^{2/3})$ remains an open problem.
Similarly, a special case of Theorem B, the implication (4), is the essentially conditional inequality denoted (${\cal I}5^{\prime}$) in [KR13]. From [KR13, theorem 3 (claim 5)] it follows that the bound in Theorem B is asymptotically tight, i.e., the term $O(\sqrt{\varepsilon})$ in the conclusion of the theorem cannot be replaced by $o(\sqrt{\varepsilon})$.
We do not know whether the term $O(\varepsilon\cdot\log\varepsilon^{-1})$ in Theorem C is asymptotically tight. This seems to be an important question, since if the bound is tight, it will imply that neither $\Gamma_{4}^{*}$ nor $\bar{\Gamma}_{4}^{*}$ are semi-algebraic sets, see [Csi14, GMM17].
Remark 2.
Theorems A and B can be derived from the infinite families of non-Shannon–type inequalities
$I(X:Y)$ $\leq I(X:Y|A)+I(X:Y|B)+I(A:B)$
$\quad{}+\frac{1}{k}I(X:Y|A)+\frac{k+1}{2}\big(I(A:X|Y)+I(A:Y|X)\big)\quad\text{for}\ k\in\mathbb{N}$
and
$I(X:Y)$ $\leq I(X:Y|A)+I(X:Y|B)+I(A:B)$
$\quad+\frac{1}{k}I(A:X|Y)+\frac{k+1}{2}\big(I(X:Y|A)+I(A:Y|X)\big)\quad\text{for }k\in\mathbb{N}$
proven by Matúš in [Mat07] (it suffices to choose a suitable value of $k$). In fact, our proof of the lemma above runs parallel to the argument of Matúš — we apply the Copy Lemma and the Ahlswede–Körner lemma, while Matúš used in [Mat07] a similar (and formally equivalent) method of adhesive extensions.
Likewise, Theorem C follows from the exponential family of inequalities in [DFZ11, Theorem 10],
$(2^{s-1}-1)I(X:Y)$ $\leq 2^{s-1}I(X:Y|A)+(2^{s-1}-1)I(X:Y|B)$
$\quad+(s-1)2^{s-2}\big(I(X:A|Y)+I(Y:A|X)\big)$
$\quad+((s-3)2^{s-2}+1)\big(I(X:B|Y)+I(Y:B|X)\big)$
$\quad+(2^{s-1}-1)I(A:B)\quad\text{for $s\in\mathbb{N}$}$
by a suitable choice of the parameter $s$.
This article presents more direct proofs of Theorems ABC without deducing preliminary infinite families of non-Shannon–type inequalities for entropy. This approach offers, in our view, several noteworthy advantages. To begin with, it is conceptually more intuitive, as it is grounded in certain structural properties of probability distributions and in the intuition of materializable common information, which we elaborate on below. This perspective enables us to reveal and prove properties that are stronger than a mere conditional form of the Ingleton inequality; see Theorem F below. In addition, the structural property we identify appears particularly amenable to computer-assisted proof systems. We believe that this may enable computer-assisted discoveries of new conditional information inequalities, in much the same spirit as non-Shannon–type inequalities and their applications have been established in [DFZ11] and especially in [Far+18, GR19].
We now outline the “structural property” of a distribution that is at the core of all our proofs. Somewhat informally, it states that if $I(A:X|Y)$ and $I(A:Y|X)$ are sufficiently small, then a part of the mutual information between $X$ and $Y$ can be “materialized” in a sense analogous to the notion of Gács–Körner common information [GK73]. This means that there exists a random variable $W$ that captures a portion of the mutual information between $X$ and $Y$: on the one hand, this $W$ has negligible conditional entropy given either $X$ or $Y$; on the other hand, $W$ retains a significant amount of information: conditioned on $W$, the mutual information between $(X,Y)$ and $A$ becomes very small. The lemma below is our main technical tool.
Lemma D (name=,restate=lWkbounds).
Suppose a tuple $(X,Y,A,B,\dots)$ is such that for some $\varepsilon>0$
$\max\left\{I(A:X|Y),I(A:Y|X)\right\}\leq\varepsilon\cdot I(X:Y)$
Then there exists a sequence of $W_{i}$, $i\in\mathbb{N}_{0}$ such that
  1. (a)
    $H(W_{i}|X,Y)=0$
  2. (b)
    $I(X:Y|W_{i+1})=I(X:Y|W_{i},A)$
  3. (c)
    $\max\left\{H(W_{i}|X),H(W_{i}|Y)\right\}\leq i\cdot\varepsilon\cdot I(X:Y)$
Moreover, if the variables $A$ and $B$ were symmetric in the original distribution $(X,Y,A,B,\ldots)$, then we may assume that this property is preserved for the extended distribution.
Corollary E (name=,restate=corWbound).
Suppose a tuple $(X,Y,A,B,\dots)$ is such that for some $\varepsilon>0$
$\max\left\{I(A:X|Y),I(A:Y|X)\right\}\leq\varepsilon\cdot I(X:Y)$
Then there exists a $W$ such that
  1. (a)
    $H(W|X,Y)=0$
  2. (b)
    $I(X,Y:A|W)=O(\sqrt{\varepsilon})\cdot I(X:Y)$
  3. (c)
    $\max\left\{H(W|X),H(W|Y)\right\}=O(\sqrt{\varepsilon})\cdot I(X:Y)$
Here an important clarification is in order. Lemma D is not literally valid for tuples of jointly distributed random variables $(X,Y,A,B,\dots)$. A version of the lemma formulated directly for random variables would necessarily be considerably more cumbersome:
“For every $\varepsilon^{\prime}>0$ for all sufficiently large $k$, for the scaling of the original distribution $(X,Y,A,B)$ obtained by taking $k$ independent, identically distributed copies of the variables, there exists a variable $W_{i}$ satisfying conditions (a), (b), and (c) within accuracy $\varepsilon^{\prime}\cdot k$.”
Although this would technically suffice, it would obscure the structural features on which our analysis relies and complicate future integration of the lemma into automated proof frameworks. For this reason, Lemma D is stated not for distributions of random variables themselves, but for tropicalized random variables, which are abstract objects derived from ordinary random variables (see [MP20]). The definition and basic properties of tropicalized variables are discussed in Section 2.2. The tropicalization procedure provides a convenient “interface” that hides certain technical steps — such as scaling up the distribution and controlling sublinear slack terms — and thereby allows us to give a cleaner and more usable formulation of the technical Lemma D and Corollary E. By contrast, our main results (Theorems ABC, and F), which rely on Lemma D and Corollary E, can be understood in the more usual way: they apply directly to conventional tuples of random variables.
We prove Lemma D and Corollary E in Section 3. The proofs may be viewed as compactly capturing the underlying construction originally due to Matúš. In particular, the proof of Corollary E requires an a priori unbounded number of applications of the Copy Lemma. This fact effectively encodes an infinite family of non-Shannon–type information inequalities, even though they never appear explicitly in the argument.
At a high level, the proofs of Theorems A and B follow a similar pattern. We first apply Corollary E, and then use standard Shannon–type inequalities on the original variables together with the auxiliary variable $W$; see the full argument in Section 4. The proofs presented in this paper are entirely conventional in the sense that we explicitly write down the necessary entropy inequalities. However, the same arguments can be generated or mechanically verified with computer assistance. In practice, one may apply Corollary E in a suitable form, take the properties (a)–(c) for the resulting auxiliary variable $W$, include all basic Shannon-type inequalities involving the variables under consideration (including $W$), and then use a linear-programming solver to deduce the desired conclusion. Since the bounds in Corollary E hold only up to $O(\sqrt{\varepsilon})$, the resulting inequality must likewise be understood up to an error term of order $O(\sqrt{\varepsilon})$.
Furthermore, we combine Lemma D with symmetry considerations and obtain two additional results. One of these results demonstrates somewhat subtler properties of tuples of random variables. If the six-tuple $(X,Y,A,B,A^{\prime},B^{\prime})$ of random variables satisfies the bounds
$\max\{I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X)\}\leq\varepsilon\cdot I(X:Y),$
then the following alternative holds: either a stronger version of the Ingleton inequality for $(X,Y,A,B)$ holds
$I(X:Y)\textrm{ is much less than }I(X:Y|A)+I(X:Y|B)+I(A:B)$
or a slightly weaker version of the Ingleton inequality for $(X,Y,A^{\prime},B^{\prime})$ holds
$I(X:Y)\ \textrm{ is not much bigger than }I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime}).$
Here is the exact statement.
Theorem F (name=,restate=thsubtler).
If the joint distribution $(X,Y,A,B,A^{\prime},B^{\prime})$ satisfies
$\max\{I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X)\}\leq\varepsilon\cdot I(X:Y)$
Then
  1. (a)
    The following alternative holds: for every threshold $0<t<1$ either we have a stronger version of Ingleton’s inequality for $(A,B,X,Y)$
    $I(X:Y)\leq I(X:Y|A)+I(X:Y|B)+I(A:B)+\big(O(\sqrt{\varepsilon})-t\big)I(X:Y)$
    or a weaker version of Ingleton’s inequality for $(A^{\prime},B^{\prime},X,Y)$
    $I(X:Y)\leq I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})+\big(O(\sqrt{\varepsilon})+t\big)I(X:Y)$
  2. (b)
    Besides, we have the following inequality for six random variables
    $\big(2-O(\sqrt{\varepsilon})\big)\cdot I(X:Y)$ $\leq I(X:Y|A)+I(X:Y|B)+I(A:B)$
    $\;\;+I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})$
Note that the assumptions in the theorem above do not involve variables $A^{\prime},B^{\prime}$ in any capacity, so that the conclusion of the theorem remains true for arbitrary $A^{\prime}$, $B^{\prime}$.
Part (b) of Theorem F probably cannot be proven with more direct approaches (without applying explicitly or implicitly infinitely many instances of the Copy Lemma). This theorem does not seem to follow from any previously known results. We prove it in Section 5.
Another result in which we combine Lemma D with symmetry considerations is Theorem C. In this theorem we assume that four conditional mutual information terms are of order $\varepsilon$. Under this assumption, we obtain a version of the Ingleton bound with an error term of order $O(\varepsilon\log\varepsilon^{-1})$ (instead of $O(\sqrt{\varepsilon})$ in the previous theorems). The proof of Theorem C is presented in Section 6.
2 Preliminaries
2.1 Notation
In what follows we use the following notation and conventions. We denote $n]:=\left\{0,1,\dots,n-1\right\}$ and $2^{[n]}$ will stand for the power set of $\left\{0,1,\dots,n-1\right\}$. For a collection of jointly distributed random variables $(X_{0},\dots,X_{n-1})$ and $J\in 2^{[n]}$, we denote by $X_{J}$ the joint random variable corresponding to the subcollection $(X_{i}\;{\bm{:}}\;i\in J)$. The entropy of a random variable $X$ (respectively, the conditional entropy of $X$ given $Y$) will be denoted $H(X)$ (and, respectively, $H(X|Y)$). We use the standard notation for the mutual information in pairs and triples of random variables,
$I(X:Y):=H(X)-H(X|Y),$
$I(X:Y|Z):=H(X|Z)-H(X|Y,Z),$
$I(X:Y:Z):=I(X:Y)-I(X:Y|Z),$
$I(X:Y:Z|W):=I(X:Y|W)-I(X:Y|W,Z).$
2.2 Tropicalization/Tensorization
In this article, we systematically use the concept of tropicalization/tensorization of random variables. This is a useful tool to investigate the asymptotic behavior of entropy profiles and to prove entropic inequalities for individual tuples of random variables.
The idea of tropicalization is the following. Given an $n$-tuple $(X_{i}:i=0,\dots,n-1)$ of random variables, we replace it with the sequence of its tensor powers $(X_{i}^{\otimes k}:i=0,\dots,n-1)_{k=1}^{\infty}$, where $X^{\otimes k}$ stands for the joint of $k$ independent copies of $X$. For such sequences we can define all the usual notions and operations, such as entropy profile, joints, etc. by performing these operations elementwise in the sequence and normalizing.
However, the class of tensor-power sequences as above is too restrictive, and we relax the definition in the following way: we consider sequences of $n$-tuples
$(\mathbf{X}^{(k)})_{k=1}^{\infty}=(X_{i}^{(k)}\;{\bm{:}}\;i=0,\dots,n-1)_{k=1}^{\infty}$
that do not deviate from tensor-powers-sequences too much, more specifically
$\operatorname{\mathsf{dist}}(\mathbf{X}^{(k)}\otimes\mathbf{X}^{(l)},\mathbf{X}^{(k+l)})\leq C\cdot\varphi(k+l)$
where $\operatorname{\mathsf{dist}}$ stands for the entropic distance between $n$-tuples of random variables and $\varphi$ is an admissible (sufficiently slowly growing) function, and $C$ is some positive constant that may depend on $\mathbf{X}$. For example, we can take $\varphi(t):=t^{3/4}$; see [MP20] for details.
We say that two such sequences $\mathbf{X}^{(k)}=(X_{i}^{(k)}\;{\bm{:}}\;i=0,\dots,n-1)$ and $\mathbf{Y}^{(k)}=(Y_{i}^{(k)}\;{\bm{:}}\;i=0,\dots,n-1)$ are equivalent if
$\lim_{k\to\infty}\frac{1}{k}\operatorname{\mathsf{dist}}(\mathbf{X}^{(k)},\mathbf{Y}^{(k)})=0$
We call equivalence classes of such sequences tropical $n$-tuples of random variables.
From now on, when we say a (tuple of) random variables we mean tropical (tuple of) random variables. Entropy of such a tropical random variable is defined as the limit of scaled entropies of individual members of the sequence.
Two observations are in order. First of all, any linear inequality that is valid for entropies for all tropical random variables implies a similar inequality for entropies of conventional random variables. Secondly, if every component of a tropical tuple obeys a given linear inequality with a sublinear error term, e.g., up to $O(1)$, then the tropical tuple itself satisfies this inequality exactly (with no error). Thus, these “tropical” objects can be interpreted as a helpful tool to study properties of the more conventional objects such as probability distributions.
2.3 Symmetries in distribution and tropical random variables
In this section we show how to transform any tropical tuple $(X_{0},\dots,X_{n-1},A,B)$ into another tuple $(X_{0},\dots,X_{n-1},\bar{A},\bar{B})$ that is symmetric with respect to the transposition of $\bar{A}$ and $\bar{B}$. This observation will allow us in some cases to assume, without loss of generality, that the distributions involved in the argument satisfy the condition of symmetry.
Lemma 3.
For every tropical tuple $(X_{0},\dots,X_{n-1},A,B)$ there exists another tropical tuple $(X_{0},\dots,X_{n-1},\bar{A},\bar{B})$ such that
  • the probability distributions in the new tuple are symmetric with respect to the transposition of $\bar{A}$ and $\bar{B}$, and
  • for any $J\in 2^{[n]}$ one has
    $H(X_{J},\bar{A})=H(X_{J},\bar{B})=\frac{1}{2}\big(H(X_{J},A)+H(X_{J},B)\big)$ and
    $H(X_{J},\bar{A},\bar{B})=H(X_{J},A,B)$
Thus, any linear combination of entropies of $(X_{0},\dots,X_{n-1},A,B)$ that is symmetric with respect to swapping $A$ and $B$ will remain unchanged when $(A,B)$ is replaced by $(\bar{A},\bar{B})$.
Proof.
Let $(X_{0},\dots,X_{n-1},A,B)$ be an $(n+2)$-tuple of random variables (not tropicalized). Denote by $\mathcal{A}$ and $\mathcal{B}$ the alphabets of $A$ and $B$, respectively, and by $p$ the joint distribution of the whole $(n+2)$-tuple. We tacitly assume that $\mathcal{A}$ and $\mathcal{B}$ are disjoint. Define two new random variables $\bar{A}$ and $\bar{B}$ jointly distributed with $X_{i}$’s, each taking values in $\mathcal{A}\cup\mathcal{B}$ and distributed according to the law $\bar{p}$ defined by
$\bar{p}(x_{0},\dots,x_{n-1},a,b)=\begin{cases}\frac{1}{2}p(x_{0},\dots,x_{n-1},a,b)&\text{if $(a,b)\in\mathcal{A}\times\mathcal{B}$}\\ \frac{1}{2}p(x_{0},\dots,x_{n-1},b,a)&\text{if $(b,a)\in\mathcal{A}\times\mathcal{B}$}\\ 0&\text{otherwise}\end{cases}$
Distribution $\bar{p}$ is symmetric with respect to swapping $\bar{A}$ and $\bar{B}$ and for any $J\in 2^{[n]}$ holds
$H(X_{J},\bar{A})=H(X_{J},\bar{B})=\frac{1}{2}(H(X_{J},A)+H(X_{J},B))+\log 2$
and
$H(X_{J},\bar{A},\bar{B})=H(X_{J},A,B)+\log 2$
Thus, we symmetrize the distribution, while changing any $(A,B)$-symmetric entropic expressions by $O(1)$, uniformly with respect to the data. Given a tropical tuple we can perform the above procedure elementwise in the sequence of tuples, thereby obtaining exact equalities for the symmetric entropic expressions. ∎
Lemma 3 implies immediately the following corollary.
Corollary 4.
Let $(X_{0},\dots,X_{n-1},A,B)$ be a tropical tuple such that for any $J\in 2^{[n]}$ one has $H(X_{J},A)=H(X_{J},B).$ Then there exists another tropical tuple $(X_{0},\dots,X_{n-1},\bar{A},\bar{B})$ such that
  • the probability distribution of the new tuple is symmetric with respect to the transposition of $\bar{A}$ and $\bar{B}$, and
  • for any $J\in 2^{[n]}$ one has
    $H(X_{J},\bar{A})=H(X_{J},A),$
    $H(X_{J},\bar{B})=H(X_{J},B),$
    $H(X_{J},\bar{A},\bar{B})=H(X_{J},A,B).$
Thus, if the entropy profile of a tuple is symmetric, we can replace it with a tuple with the same entropy profile and a symmetric distribution.
Remark 5.
Let ${\cal L}_{i}(X,Y,A,B)$, $i=1,2$, be linear combinations of entropies involving $X,Y,A,B$. Assume also that these ${\cal L}_{i}$ are symmetric (as formal expressions) with respect to the transposition of the symbols $A$ and $B$, e.g. the expression from the Ingleton inequality:
$\begin{array}[]{l}I(A:B|X)+I(A:B|Y)+I(X:Y)-I(A:B)\\ =H(A,X)+H(B,X)-H(A,B,X)+H(A,Y)+H(B,Y)-H(A,B,Y)\\ \phantom{=}-H(X,Y)-H(A)-H(B)+H(A,B)\end{array}$
In this setting, if we can prove a statement of the form
${\cal L}_{1}(X,Y,A,B)=0\Rightarrow{\cal L}_{2}(X,Y,A,B)\geq 0$
for the distributions symmetric with respect to the transposition of $A$ and $B$, then this implication is true for all distributions. Indeed, given an arbitrary distribution $(X,Y,A,B)$, we can apply Lemma 3 and construct its symmetrized version. If ${\cal L}_{1}(X,Y,A,B)=0$ for the original distribution, the same equality holds for the symmetrized one (here we use symmetry of ${\cal L}_{1}$). So we can conclude that ${\cal L}_{2}(X,Y,A,B)\geq 0$ for the symmetrized distribution. Finally, we use symmetry of ${\cal L}_{2}$ and conclude that the same inequality holds for the original distribution.
This observation allows us in some cases to assume without loss of generality that the random variables under consideration have the property of symmetry.
2.4 Copy lemma
In this section, we recall the classical Copy Lemma. Essentially, it states that for every distribution $(X,Y,A)$, one can construct an extended distribution $(X,Y,A,A^{\prime})$ such that the conditional distribution of $A^{\prime}$ given $X$ coincides with that of $A$ given $X$, and moreover, the random variables form the Markov chain
$(Y,A)\to X\to A^{\prime}.$
In other words, $A^{\prime}$ is an “independent copy” of $A$ that preserves the same joint distribution with $X$. The lemma extends to the setting where $X$ and $Y$ are not individual random variables but jointly distributed tuples.
Despite its simplicity, this lemma remains a powerful tool in proving non-Shannon–type information inequalities for entropy. This construction appeared implicitly in the original work of Zhang and Yeung [ZY97, ZY98], and was later formulated explicitly in [DFZ06].
Lemma 6 (Copy Lemma).
Every tuple of jointly distributed random variables
$(X_{i},Y_{j},A\;{\bm{:}}\;i=0,\ldots,k-1;\;j=0,\dots,\ell-1)$
can be extended to a longer tuple
$(X_{i},Y_{j},A^{\prime}\;{\bm{:}}\;i=0,\ldots,k-1;\;j=0,\dots,\ell-1)$
such that $(X_{0},\ldots,X_{k-1},A)$ and $(X_{0},\ldots,X_{k-1},A^{\prime})$ have the same distribution, and $A^{\prime}$ is independent of $(Y_{0},\ldots,Y_{\ell-1},A)$ conditioned on $(X_{0},\ldots,X_{k-1})$.
Moreover, if the original distribution of $(X_{[k]},Y_{[\ell]},A)$ was symmetric with respect to some permutation of variables $(Y_{j},A)$ and/or permutation of variables $(X_{i})$, then the extended distribution of $(X_{[k]},Y_{[\ell]},A,A^{\prime})$ will possess the same symmetries.
Proof.
The new variable $A^{\prime}$ will have the same alphabet as $A$, and the new probability distribution can be defined by the following equation: for all instances $a$ and $a^{\prime}$ of $A$ and $A^{\prime}$, respectively,
$\mathrm{Pr}[(X_{[k]},$ $Y_{[\ell]},A,A^{\prime})=(\mathbf{x},\mathbf{y},a,a^{\prime})]=$
$=\begin{cases}\frac{\mathrm{Pr}[(X_{[k]},Y_{[\ell]},A)=(\mathbf{x},\mathbf{y},a)]\cdot\mathrm{Pr}[(X_{[k]},A)=(\mathbf{x},a^{\prime})]}{\mathrm{Pr}[X_{[k]}=\mathbf{x}]}&\text{if $\mathrm{Pr}[X_{[k]}=\mathbf{x}]>0$,}\\ 0&\text{otherwise}.\end{cases}$
where $\mathbf{x}=(x_{0},\dots,x_{k-1})$, $\mathbf{y}=(y_{0},\dots,y_{\ell-1})$ are instances of $X_{[k]}$ and $Y_{[\ell]}$, respectively.
The moreover claim of the lemma follows directly from the definition of the new distribution. ∎
The entropy profile of the extended tuple $(X_{[k]},Y_{[\ell]},A,A^{\prime})$ is not totally defined by the entropies of the original distribution $(X_{[k]},Y_{[\ell]},A)$. However, some components of the new entropy profile are determined by the definition of $A^{\prime}$: for all $I\subset[k]$ and $J\subset[\ell]$
$H(X_{I},A^{\prime})$ $=H(X_{I},A),$
$H(Y_{J},A,A^{\prime}\mid X_{[k]})$ $=H(Y_{J},A\mid X_{[k]})+H(A^{\prime}\mid X_{[k]}).$
2.5 Ahlswede–Körner reduction
The Ahlswede–Körner reduction, [AK06, AK77], modifies a given tropical $n$-tuple so that some entropies in the new tuple vanish, while some (combinations of) other entropies stay the same.
Lemma 7 (Ahlswede–Körner reduction).
Let us consider a tropical $n$-tuple $(X_{i}:i=0,\dots,n-1)$ and a subcollection $(X_{i}:i=0,\dots,k-1)$ of it, where $k\leq n$. Then there exists an $(n+1)$-tuple $(X_{0},\dots,X_{n-1},X^{\prime}_{n-1})$, such that
$\begin{cases}H(X^{\prime}_{n-1}|X_{0},\dots,X_{k-1})=0\\ H(X_{J}|X^{\prime}_{n-1})=H(X_{J}|X_{n-1})&\text{for any $J\in 2^{[k]}$}\end{cases}$
We will call the new random variable $X^{\prime}_{n-1}$ the Ahlswede–Körner reduction of $X_{n-1}$ under $(X_{0},\dots,X_{k-1})$ and denote the new variable
$X^{\prime}_{n-1}:=\operatorname{\mathsf{AK}}(X_{n-1}\mid X_{0},\dots,X_{k-1}).$
3 Main construction
In this section we prove our main technical results, Lemma D and Corollary E, which are repeated below verbatim. They are systematically applied in all proofs of the article. Assume we are given a jointly distributed triple $(X,Y,A)$ or, possibly, a longer joint distribution $(X,Y,A,B,C_{0},\ldots,C_{n-1})$ extending $(X,Y,A)$. For each integer $k\geq 0$, we extend this tuple with a new variable $W_{k}$ satisfying the properties stated in the lemma.
[…]
3.1 The main construction
We start with a tuple of random variables $(X,Y,A,B,C_{0},\dots,C_{n-1})$, possibly satisfying some symmetry that permutes variables $A$, $B$ and $C_{i}$’s, and then apply the following steps.
Step 1 (duplication):
First of all, we apply the Copy Lemma $k$ times to $A$ under $(X,Y)$ and obtain new variables $A^{\prime}_{1},\ldots,A^{\prime}_{k}$ such that
  • for each $i=1,\ldots,k$ the conditional distribution of $A^{\prime}_{i}$ given $(X,Y)$ is the same as the conditional distribution of $A$;
  • the random variables $A^{\prime}_{1},\ldots,A^{\prime}_{k}$ and the tuple $(A,B,C_{0},\ldots,C_{n-1})$ are all together mutually independent conditional on $(X,Y)$.
Denote $\tilde{A}_{k}=(A^{\prime}_{1},\ldots,A^{\prime}_{k})$. Observe that if the original distribution was symmetric under swapping $A$ and $B$, then the same symmetry holds for the extended distribution
$(X,Y,A,B,C_{0},\ldots,C_{n-1},\tilde{A}_{k})$
(see the moreover claim of the Copy Lemma).
Step 2 (reduction):
Further, we apply the Ahlswede–Körner reduction to $\tilde{A}_{k}$ under $(A,B,C_{0},\ldots,C_{n-1})$ and denote the result by
$W_{k}:=\operatorname{\mathsf{AK}}(\tilde{A}_{k}\mid X,Y,A,B,C_{0},\ldots,C_{n-1}).$
We use the convention that $W_{0}$ is the trivial (valued in a one-point-set) random variable.
[…]
$H(A,W_{k})=H(B,W_{k}),\quad H(A,C_{J},W_{k})=H(B,C_{J},W_{k}),\text{ etc}.$
3.2 Proof of Lemma D
We use $W_{[k]}$ constructed in the previous section to prove the lemma. To prove assertion (a), we note that by the Copy Lemma, the variable $\tilde{A}_{k}$ is independent of $(A,B,C_{0},\ldots,C_{n-1})$ conditioned on $(X,Y)$. The same holds for $W_{k}=\operatorname{\mathsf{AK}}(\tilde{A}_{k}|X,Y,A,B,C_{0},\dots,C_{n-1})$, by the definition of the Ahlswede–Körner reduction. Thus we have
$H(W_{k}|X,Y)=H(W_{k}|X,Y,A,B,C_{0},\dots,C_{n-1})=0$
To prove assertion (b) it suffices to observe that the conditional distributions of $(X,Y)$ given $(A^{\prime}_{1},\ldots,A^{\prime}_{k+1})$ and given $(A^{\prime}_{1},\ldots,A^{\prime}_{k},A)$ are identical.
To prove (c), we observe that
$H(W_{k}|X)$ $=I(\tilde{A}_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1}|X)$
/* from the AK Lemma */
$=I(\tilde{A}_{k}:X,Y|X)$
/* since $\tilde{A}_{k}\to(XY)\to(A,B,C_{0},\ldots,C_{n-1})$ is a Markov chain */
$=I(\tilde{A}_{k}:Y|X)=H(\tilde{A}_{k}|X)-H(\tilde{A}_{k}|X,Y).$
Let us estimate the terms in the right-hand side of this equality. On the one hand,
$H(\tilde{A}_{k}|X)=H(A^{\prime}_{1},\ldots,A^{\prime}_{k}|X)\leq\sum\limits_{i=1}^{k}H(A^{\prime}_{i}|X)=k\cdot H(A|X).$
On the other hand, by the definition of $A^{\prime}_{i}$, we have
$H(\tilde{A}_{k}|X,Y)=H(A^{\prime}_{1},\ldots,A^{\prime}_{k}|X,Y)=\sum\limits_{i=1}^{k}H(A^{\prime}_{i}|X,Y)=k\cdot H(A|X,Y).$
Thus, $H(W_{k}|X)=H(\tilde{A}_{k}|X)-H(\tilde{A}_{k}|X,Y)$ rewrites to
$H(W_{k}|X)=k\cdot\left(H(A|X)-H(A|X,Y)\right)=k\cdot I(A:Y|X)\leq k\cdot\varepsilon\cdot I(X:Y).$
A similar argument gives the bound for $H(W_{k}|Y)$.
It remains to prove the moreover part of the lemma. We already know that the entropy profile $(X,Y,A,B,C_{0},\ldots,C_{n-1},W_{k})$ is symmetric under swapping $A$ and $B$. To make the property stronger and enforce the symmetry of the tropical tuple, we apply Corollary 4.
3.3 Proof of Corollary E
[…]
Proof.
The argument is based on Lemma D, where we will choose a suitable $k$ and let $W=W_{k}$. We apply Lemma D with the minimal $k$ such that
$I(X:Y|W_{k})\leq I(X:Y|A,W_{k})+\delta\cdot I(X:Y)$
(the choice of the parameter $\delta<1$ is explained later). This means that for each $i=1,2,\ldots,k$ the gap
$I(X:Y|W_{i-1})-I(X:Y|W_{i})$
is at least $\delta\cdot I(X:Y)$. Since the sum of all such gaps does not exceed $I(X:Y)$, the number of steps $k$ is at most $k\leq 1/\delta.$ From Lemma D(c) we get
$\max\left\{H(W_{k}|X),H(W_{k}|Y)\right\}\leq k\cdot\varepsilon\cdot I(X:Y)\leq\frac{\varepsilon}{\delta}\cdot I(X:Y)$
It follows that
$I($ $A:X,Y|W_{k})=I(A:X|Y,W_{k})+I(A:Y|X,W_{k})+I(A:X:Y|W_{k})$
$\leq I(A:X|Y)+H(W_{k}|X)+I(A:Y|X)+H(W_{k}|Y)+I(A:X:Y|W_{k})$
$\leq I(A:X|Y)+I(A:Y|X)+I(A:X:Y|W_{k})+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
$\leq I(A:X:Y|W_{k})+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
$\leq I(X:Y|W_{k})-I(X:Y|A,W_{k})+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
$\leq\delta\cdot I(X:Y)+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
We set $\delta=\sqrt{\varepsilon}$, and the corollary follows. ∎
4 Conditional inequalities with precision $O(\sqrt{\varepsilon})$
In this section we prove that if the conditional information quantities
I(A:X|Y),I(A:Y|X)I(A:X|Y),\ I(A:Y|X)
are of order $\varepsilon$, then a version of the Ingleton inequality (ING) is valid with a small correction term of order $\sqrt{\varepsilon}$; see Theorem A.
[…]
Proof.
First of all, we apply Corollary E and take a $W$ such that
  • $H(W|X)=O(\sqrt{\varepsilon})\cdot I(X:Y)$,
  • $H(W|Y)=O(\sqrt{\varepsilon})\cdot I(X:Y)$,
  • $I(A:X,Y|W)=O(\sqrt{\varepsilon})\cdot I(X:Y)$.
The rest of the proof is an application of suitable Shannon–type inequalities and can therefore be verified with the help of any solver that handles Shannon–type entropy inequalities.
For all tuples $(X,Y,A,W)$ the following Shannon–type inequality holds:
$I(X:Y|A)\geq H(W|A)+I(X:Y|A,W)-\big(H(W|X)+H(W|Y)\big),$
see inequality (24) in the Appendix. For the $W$ chosen above, this inequality implies
$\begin{split}I(X:Y|A)&\geq H(W|A)+I(X:Y|A,W)-O(\sqrt{\varepsilon})\cdot I(X:Y)\\ &\geq H(W|A)+I(X:Y|W)-O(\sqrt{\varepsilon})\cdot I(X:Y).\end{split}$
It follows that
$\begin{split}H(W|A)&\leq I(X:Y|A)-I(X:Y|W)+O(\sqrt{\varepsilon})\cdot I(X:Y)\end{split}$ (5)
Now we adapt the conventional inference of the Ingleton inequality. We use the following Shannon–type equalities (see (21), (22), (23) in the Appendix):
$\left.\begin{aligned} H(W|B)&\leq H(W|X)+H(W|Y)+I(X:Y|B)\\ &\leq I(X:Y|B)+O(\sqrt{\varepsilon})\cdot I(X:Y|A)\\ &(\text{follows from the choice of $W$)},\\ H(W)&\leq H(W|A)+H(W|B)+I(A:B),\\ I(X:Y)&\leq H(W)+I(X:Y|W)\end{aligned}\right\}$ (6)
Summing inequalities (5) and (6) yields the conclusion of the theorem. ∎
Remark 8.
If the quantity $I(X:Y)$ is a priori bounded, e.g., if $I(X:Y)\leq 1$, then the inequality in Theorem A rewrites to
$I(X:Y)\leq I(X:Y|A)+I(X:Y|B)+I(A:B)+O(\sqrt{\varepsilon}).$
[…]
Remark 9.
Because of the bound $I(X:Y|A)\leq\varepsilon\cdot I(X:A)$ in the assumptions, we can equivalently rewrite the conclusion of Theorem B as
$I(X:Y)\leq I(X:Y|B)+I(A:B)+O(\sqrt{\varepsilon})\cdot I(X:A).$
Proof.
Similarly to the previous argument, we apply Corollary E (switching $A$ and $Y$ in the assumptions and the conclusion of the lemma) and take a $W$ such that
  • $H(W|X)=O(\sqrt{\varepsilon})\cdot I(X:A)$,
  • $H(W|A)=O(\sqrt{\varepsilon})\cdot I(X:A)$,
  • $I(X,A:Y|W)=O(\sqrt{\varepsilon})\cdot I(X:A)$.
For brevity, in what follows we use the notation $\Delta:=\sqrt{\varepsilon}\cdot I(X:A)$. The rest of this proof is an application of suitable Shannon–type inequalities.
We begin with a Shannon–type inequality
$H(W)\leq H(W|A)+H(W|B)+I(A:B),$
see (21) in the Appendix. The term $H(W|A)$ in the right-hand side of the inequality above is bounded by $O(\Delta)$, so we can rewrite this inequality as
$H(W)\leq H(W|B)+I(A:B)+O(\Delta).$ (7)
To bound the term $H(W|B)$, we use another Shannon–type inequality
$H(W|B)\leq H(W|X)+H(W|Y)+I(X:Y|B),$
see (22) in the Appendix. The term $H(W|X)$ in the right-hand side of the last inequality can be replaced immediately by $O(\Delta)$, and we obtain
$H(W|B)\leq H(W|Y)+I(X:Y|B)+O(\Delta).$ (8)
Our next step is an upper bound for $H(W|Y)$. We isolate this step of the proof into a separate claim.
Claim 10.
$H(W|Y)\leq H(W)-I(X:Y)+O(\Delta).$
Proof of the claim.
First, applying several times the chain rule for the mutual information, we obtain
$\begin{split}I(X:Y)&\leq I(X:Y)+I(W:Y|X)\\ &=I(X,W:Y)\\ &=I(W:Y)+I(X:Y|W),\\ &\leq I(W:Y)+I(X:Y|W)+I(A:Y|X,W)\\ &=I(W:Y)+I(X,A:Y|W)\\ &=I(W:Y)+O(\Delta),\end{split}$
where the last step uses the assumption $I(X,A:Y|W)=O(\Delta)$.
It follows that
[…]
and the claim is proven. ∎
Combining inequalities (7), (8), and the Claim, we obtain
$H(W)\leq H(W)-I(X:Y)+I(X:Y|B)+I(A:B)+O(\Delta),$
By Remark 9 this completes the proof of the theorem. ∎
5 Subtler claims with the $O(\sqrt{\varepsilon})$ precision
In this section we combine Lemma D with symmetry considerations, derive a new information inequality for six-tuples of random variables, and prove Theorem F from the introduction.
[…]
Proof.
First of all, we observe that the inequalities in both claims (a) and (b) are symmetric as syntactic expressions with respect to the transposition of $A$ and $B$. Therefore, we can apply Lemma 3 (see also Remark 5) and assume without loss of generality that $A$ and $B$ are symmetric in the distribution $(X,Y,A,B,A^{\prime},B^{\prime})$. Now we are ready to prove the theorem. To address the first claim of the theorem, we apply Corollary E to construct $W$ such that
  • $H(W|X)=O(\sqrt{\varepsilon})\cdot I(X:Y)$,
  • $H(W|Y)=O(\sqrt{\varepsilon})\cdot I(X:Y)$,
  • $I(A:X,Y|W)=O(\sqrt{\varepsilon})\cdot I(X:Y)$.
Due to Lemma D (the moreover part) we may assume that all entropy quantities for $(X,Y,A,W)$ are the same as the corresponding quantities for $(X,Y,B,W)$. Similarly to (5), using the symmetry we have
$H(W|A)$ $\leq I(X:Y|A)-I(X:Y|W)+O(\sqrt{\varepsilon})\cdot I(X:Y)$
$H(W|B)$ $\leq I(X:Y|B)-I(X:Y|W)+O(\sqrt{\varepsilon})\cdot I(X:Y)$
The sum of two inequalities above together with Shannon–type inequalities
$I(X:Y)$ $\leq H(W)+I(X:Y|W),$
$H(W)$ $\leq H(W|A)+H(W|B)+I(A:B)$
gives
$\begin{split}I(X:Y)&\leq I(X:Y|A)+I(X:Y|B)+I(A:B)-I(X:Y|W)\\ &+O(\sqrt{\varepsilon})\cdot I(X:Y)\end{split}$ (9)
Therefore, if $t\leq I(X:Y|W)/I(X:Y)$, then
$I(X:Y)\leq I(X:Y|A)+I(X:Y|B)+I(A:B)+\big(O(\sqrt{\varepsilon})-t)\cdot I(X:Y).$
This gives the first part of the alternative in claim (a).
On the other hand, for all $(X,Y,A^{\prime},B^{\prime},W)$ we can follow the usual scheme of the proof of the Ingleton’s inequality with $W$ representing the common information of $X$ and $Y$. More formally, we add Shannon–type inequalities
$H(W)$ $\leq H(W|A^{\prime})+H(W|B^{\prime})+I(A^{\prime}:B^{\prime}),$
$H(W|A^{\prime})$ $\leq H(W|X)+H(W|Y)+I(X:Y|A^{\prime}),$
$H(W|B^{\prime})$ $\leq H(W|X)+H(W|Y)+I(X:Y|B^{\prime}),$
$I(X:Y)$ $\leq H(W)+I(X:Y|W)$
to obtain
$I(X:Y)$ $\leq I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})$
$+2\big(H(W|X)+H(W|Y)\big)+I(X:Y|W)$
Therefore, for the chosen $W$ we have
$\begin{split}I(X:Y)&\leq I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})\\ &+O(\sqrt{\varepsilon})\cdot I(X:Y)+I(X:Y|W)\end{split}$ (10)
Thus, if $t\geq I(X:Y|W)/I(X:Y)$, we obtain
$I(X:Y)\leq I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})+\big(O(\sqrt{\varepsilon})+t\big)\cdot I(X:Y).$
This gives the second part of the alternative.
To prove part (b) of the theorem, we take the sum of inequalities (9) and (10). ∎
6 Inequalities with the precision of $O(\varepsilon\log\frac{1}{\varepsilon})$
This section is devoted to the proof of Theorem C. For convenience the statement is repeated below verbatim.
[…]
Proof.
Since the inequalities in both the assumption and the conclusion of the theorem are symmetric with respect to swapping $A$ and $B$, we may apply Lemma 3 to the four-tuple $(X,Y,A,B)$. Thus, without loss of generality, we may assume that the distribution is symmetric (see also Remark 5).
Now we use Lemma D and symmetry of the distribution to find a sequence of extensions $(X,Y,A,B,W_{i})$, $i=0,1,\dots$ satisfying
$H(W_{i}|X,Y)$ $=0$
$I(X:Y|W_{i+1})$ $=I(X:Y|W_{i},A)$
$I(X:Y|W_{i+1})$ $=I(X:Y|W_{i},B)$
$\max\left\{H(W_{i}|X),H(W_{i}|Y)\right\}$ $\leq i\cdot\varepsilon\cdot I(X:Y)$
Fix two numbers $\lambda,\delta\in(0,1)$. We postpone the concrete choices for these two parameters for later in the proof. Take the largest $k$ such that for all $0\leq i<k$ holds
$I(X:Y|W_{i+1})$ $\leq\lambda\cdot I(X:Y|W_{i})\quad\text{and}$ (11)
$I(X:Y|W_{i})$ $\geq\delta\cdot I(X:Y)$ (12)
Then
$\delta\cdot I(X:Y)\leq I(X:Y|W_{k})\leq\lambda I(X:Y|W_{k-1})\leq\ldots\leq\lambda^{k}\cdot I(X:Y)$
Therefore
$k\leq\frac{\ln\delta}{\ln\lambda}$ (13)
There are two possible cases, depending on whether condition (11) or (12) is violated at $i=k$.
6.1 Case 1: $I(X:Y|W_{k+1})\geq\lambda\cdot I(X:Y|W_{k})$
We use inequality (24) from the Appendix to obtain
$\begin{split}I(X:Y|A)&\geq H(W_{k}|A)+I(X:Y|W_{k},A)-\big(H(W_{k}|X,A)+H(W_{k}|Y,A)\big)\\ &\geq H(W_{k}|A)+I(X:Y|W_{k+1})-\big(H(W_{k}|X)+H(W_{k}|Y)\big)\\ &\geq H(W_{k}|A)+\lambda\cdot I(X:Y|W_{k})-2k\cdot\varepsilon\cdot I(X:Y)\end{split}$
Similarly we have
$I(X:Y|B)\geq H(W_{k}|B)+\lambda\cdot I(X:Y|W_{k})-2k\cdot\varepsilon\cdot I(X:Y)$
We can rewrite these inequalities as
$H(W_{k}|A)$ $\leq I(X:Y|A)-\lambda\cdot I(X:Y|W_{k})+2k\cdot\varepsilon\cdot I(X:Y)$ (14)
$H(W_{k}|B)$ $\leq I(X:Y|B)-\lambda\cdot I(X:Y|W_{k})+2k\cdot\varepsilon\cdot I(X:Y)$ (15)
Applying two Shannon inequalities (21) and (23) from the Appendix we get
$H(W_{k})$ $\leq H(W_{k}|A)+H(W_{k}|B)+I(A:B)$ (16)
$I(X:Y)$ $\leq H(W_{k})+I(X:Y|W_{k})$ (17)
Adding inequalities (1417) we get
$\begin{split}I(X:Y)&\leq I(X:Y|A)+I(X:Y|B)+I(A:B)\\ &\quad+(1-2\lambda)I(X:Y|W_{k})+4k\cdot\varepsilon\cdot I(X:Y)\end{split}$ (18)
Choosing $\lambda=1/2$ and using the bound (13) for $k$ we obtain
$\begin{split}I(X:Y)&\leq I(X:Y|A)+I(X:Y|B)+I(A:B)\\ &\quad+4\varepsilon\log_{2}\delta^{-1}\cdot I(X:Y)\end{split}$ (19)
6.2 Case 2: $I(X:Y|W_{k})\leq\delta\cdot I(X:Y)$
We apply inequality (25) from the Appendix, and use the properties of $W_{i}$ from Lemma D to estimate
$I(X:Y)$ $\leq I(X:Y|A)+I(X:Y|B)+I(A:B)$
$\quad+I(X:Y|W_{k})+I(X:W_{k}|Y)+I(W_{k}:Y|X)$
$\leq I(X:Y|A)+I(X:Y|B)+I(A:B)+(\delta+2k\cdot\varepsilon)I(X:Y)$
Using (13) we rewrite the last inequality as
$\begin{split}I(X:Y)&\leq I(X:Y|A)+I(X:Y|B)+I(A:B)\\ &\quad+(\delta+2\varepsilon\log_{2}\delta^{-1})I(X:Y)\end{split}$ (20)
Now we set $\delta=\varepsilon$ in inequalities (19) and (20) and take the maximum of the two bounds. This gives
$I(X:Y)\leq I(X:Y|A)+I(X:Y|B)+I(A:B)+O(\varepsilon\log\varepsilon^{-1})I(X:Y)$
Appendix: Useful information inequalities
In this section we collect several equalities and inequalities for the Shannon entropy that are used throughout the paper. First of all, we freely use the chain rule for the mutual information
$I(X,Y:A)=I(X:A)+I(Y:A|X)$
and its “conditioned” version
$I(X,Y:A|B)=I(X:A|B)+I(Y:A|X,B).$
Besides, we apply several times the well-known Shannon–type inequality
$H(Z)\leq H(Z|X)+H(Z|Y)+I(X:Y)$ (21)
(which is equivalent to the sum of $I(X:Y|Z)\geq 0$ and $H(Z|X,Y)\geq 0$) and its conditioned version
$H(Z|W)\leq H(Z|X,W)+H(Z|Y,W)+I(X:Y|W),$ (22)
In addition, we use three other Shannon–type inequalities:
$I(X:Y)\leq H(W)+I(X:Y|W)$ (23)
(which is equivalent to the sum of $H(W|X)\geq 0$ and $I(W:X|Y)\geq 0$),
$H(W)+I(X:Y|W)\leq I(X:Y)+\big(H(W|X)+H(W|Y)\big)$
(which easily rewrites to the trivial $H(W|X,Y)\geq 0$), and its conditioned version
$H(W|A)+I(X:Y|W,A)\leq I(X:Y|A)+\big(H(W|X,A)+H(W|Y,A)\big).$ (24)
Finally, we use a version of the classical non-Shannon–type inequality from [ZY98] in the form with five variables (proven in [Mak+02]),
$\begin{split}I(X:Y)&\leq I(X:Y|A)+I(X:Y|B)+I(A:B)\\ &\quad+I(X:Y|W)+I(X:W|Y)+I(Y:W|X).\end{split}$ (25)

### v2

@@
 1 Introduction
-In this paper we study inequalities for Shannon entropy, and, in particular, variants of the Ingleton inequality in the entropic setting.
-It is well known that the Ingleton inequality in its original form does not hold for entropies of certain tuples of jointly distributed random variables; see, e.g.
-[Mat95, Csi96, ZY98, Mat99, Ham+00].
+In this paper we study inequalities for Shannon entropy, in particular variants of the Ingleton inequality in the entropic setting.
+At a high level, the goals of this paper can be summarized as two broad objectives:
+  1. (I)
+    gaining a better understanding of the geometry of the boundary of the cone of almost-entropic vectors;
+  2. (II)
+    developing tools for further progress in understanding conditional independence structures.
+More specifically, we focus on conditional versions of the Ingleton inequality and on related structural properties of entropy.
+It is well known that the Ingleton inequality in its original form does not hold for the entropies of certain collections of jointly distributed random variables; see, e.g.
+[Mat95, Csi96, ZY98, Mat99, HRSV00].
 However, under certain restrictions (or with suitable additional terms) the inequality may still hold.
-We discuss several manifestations of this important phenomenon.
+We study several manifestations of this phenomenon.
+1.1 The cone of almost-entropic points
 We begin with a brief review of the standard framework for information inequalities.
-For a tuple of random variables $(X_{i}\;{\bm{:}}\;i=0,\dots,n-1)$ its entropy profile is the function on the power set of the set $\left\{0,\dots,n-1\right\}$ indexing the tuple.
-The value of this function on $J\in 2^{[n]}$ is the entropy of the joint distribution $(X_{i}\;{\bm{:}}\;i\in J)$.
+For a tuple of random variables $(X_{i}\;{\bm{:}}\;i=0,\dots,n-1)$ its entropy profile is the function on the power set of the set $\left\{0,\dots,n-1\right\}$ indexing the tuple whose value on a set $J\in 2^{[n]}$ is the entropy of the joint distribution $(X_{i}\;{\bm{:}}\;i\in J)$.
 Since the entropy of the empty collection is always zero, we drop the corresponding value from consideration.
@@
 The set of all entropy profiles for all $n$-tuples of random variable is usually denoted $\Gamma_{n}^{\star}$.
-The points in the closure of the set of entropy profiles are called almost entropic points.
+The points in the closure of the set of entropy profiles are called almost-entropic points.
 This set is denoted by $\bar{\Gamma}_{n}^{\star}$.
@@
 The geometry of this cone is quite non-trivial.
-It is known, for example, that for all $n\geq 4$ the set $\bar{\Gamma}_{n}^{\star}$ is not polyhedral, [Mat07]).
 An exact description of this cone for $n\geq 4$ is still unknown.
-However, there are some known lower and upper bounds for this set.
+However, there are some known inner and outer bounds for this set.
 In what follows we briefly discuss the basic facts about $\bar{\Gamma}_{4}^{\star}$.
@@
 The set of all $\varphi$ satisfying these inequalities is traditionally denoted $\Gamma_{4}$ (the cone of submodular functions on a four-point set).
-This set is clearly a superset (i.e., an upper bound) of $\bar{\Gamma}_{4}^{\star}$.
-Some sufficient conditions for being an entropic point can be formulated with the help of the Ingleton inequalities. In what follows we use the notation: for a set function $\varphi$ define
+This set is clearly a superset (i.e., an outer bound) of $\bar{\Gamma}_{4}^{\star}$.
+In other words, the axioms of polymatroids (non-negativity, monotonicity, and submodularity) are necessary conditions for a point to be entropic.
+The cone $\bar{\Gamma}_{4}^{\star}$ is strictly smaller than $\Gamma_{4}$; that is, besides the inequalities valid for all polymatroids, there exist other linear inequalities that are valid for all almost-entropic points (for distributions of $n\geq 4$ random variables).
+The inequalities that are valid for all almost-entropic points but do not follow from the Shannon inequalities are called non-Shannon type inequalities.
+The first example of such an inequality was discovered by Z.
+Zhang and R.
+W.
+Yeung in [ZY98].
+In the following years many other non-Shannon type inequalities were discovered; see, e.g., [MMRV02, DFZ06, XWS08, DFZ11, Csi16].
+Moreover, F.
+Matúš proved in [Mat07b] that there exist infinitely many independent non-Shannon type inequalities (in other words, $\bar{\Gamma}_{4}^{\star}$ is not polyhedral, as well as all $\bar{\Gamma}_{n}^{\star}$ for $n>4$).
+In the early years after the discovery of the first non-Shannon type inequalities, much effort was devoted to understanding how large the gap between $\bar{\Gamma}_{4}^{\star}$ and $\Gamma_{4}$ is.
+Accordingly, particular interest was placed on non-Shannon inequalities that exclude a significant region of the cone $\Gamma_{4}$.
+In recent years, attention has shifted to somewhat different questions.
+For example: what is the geometry of the boundary of the cone $\Gamma_{4}$?
+How accurately can it be approximated by piecewise algebraic surfaces?
+How smoothly does the curved part of the boundary approach a given flat face?
+The main results of this paper are primarily relevant to the second type of questions, namely those related to the geometry of the boundary of $\bar{\Gamma}_{4}^{\star}$ (see Objective (I) on p.
+(I)).
+The main results of this paper are centered around the Ingleton inequality described in [Ing71]. To discuss this inequality, we introduce the following notation: for a set function $\varphi$ define
 $\varphi(x:y)$ $:=\varphi(\left\{x\right\})+\varphi(\left\{y\right\})-\varphi(\left\{x,y\right\}),$
 $\varphi(x:y|z)$ $:=\varphi(\left\{x,z\right\})+\varphi(\left\{y,z\right\})-\varphi(\left\{x,y,z\right\})-\varphi(\left\{z\right\}).$
-With this notation, the Ingleton inequality can be written in the following compact form:
+Now the Ingleton inequality can be written in the following compact form:
 $\varphi(0:1)\leq\varphi(0:1|2)+\varphi(0:1|3)+\varphi(2:3).$ (1)
 Permuting the set $\left\{0,1,2,3\right\}$ we obtain $\binom{4}{2}=6$ similar inequalities that differ in the order of the variables.
-We denote by $\Gamma^{\rm Ing}_{4}$ the cone of all non-negative, monotone, submodular functions $\varphi$ on $\left\{0,1,2,3\right\}$ that satisfy, in addition to the Shannon inequalities, the six versions of the Ingleton inequality on a four-point set.
-This set is a subset (lower bound) of the set of almost entropic points: it is known that every $\varphi$ in $\Gamma^{\rm Ing}_{4}$ is almost entropic, see [MS95] (independently, this was proved in [Ham+00]).
-The sets $\Gamma_{4}$ and $\Gamma_{4}^{\mathrm{Ing}}$ are both polyhedral cones, and the set of entropic points is sandwiched between the two: $\Gamma_{4}^{\mathrm{Ing}}\subsetneq\bar{\Gamma}_{4}^{\star}\subsetneq\Gamma_{4}.$ The closure of $\Gamma_{4}\setminus\Gamma_{4}^{\mathrm{Ing}}$ is the union of six closed convex cones with disjoint interiors.
+We denote by $\Gamma^{\rm Ing}_{4}$ the cone of all non-negative, monotone, submodular functions $\varphi$ on $\left\{0,1,2,3\right\}$ that satisfy, in addition to the Shannon inequalities, the six versions of the Ingleton inequality on a four-point set (this cone under a different notation was discussed, e.g., in [CGK10, WC17]).
+The Ingleton inequality is a sufficient condition for being an entropic point, i.e., every $\varphi$ in $\Gamma^{\rm Ing}_{4}$ is entropic, see [MS95] (independently, this was proved in [HRSV00]).
+In other words, $\Gamma^{\rm Ing}_{4}$ is a subset (inner bound) of the set of almost-entropic points.
+It is well known that the Ingleton inequality is not a necessary condition for being an entropic point, i.e., there exist entropic $\varphi$ outside $\Gamma^{\rm Ing}_{4}$, see [Mat95, Csi96, ZY98, Mat99, HRSV00, BN20, MC16].
+The sets $\Gamma_{4}$ and $\Gamma_{4}^{\mathrm{Ing}}$ are both polyhedral cones, and the set of almost-entropic points is sandwiched between the two: $\Gamma_{4}^{\mathrm{Ing}}\subsetneq\bar{\Gamma}_{4}^{\star}\subsetneq\Gamma_{4}.$ The closure of $\Gamma_{4}\setminus\Gamma_{4}^{\mathrm{Ing}}$ consists of six closed convex cones with disjoint interiors, [MP20].
 They are symmetric under permutations of the set $\left\{0,1,2,3\right\}$.
-From now on we consider just one of these cones in which the strict version of inequality (1) is violated, and we refer to it as the non-Ingleton cone.
+From now on, we consider one of these cones in which the strict inequality (1) is violated and, following [MP20], refer to it as the non-Ingleton cone.
+Let us briefly recall how this object can be viewed in geometric terms (see [MP20] or [CGK10, GCG11] for more details).
 The non-Ingleton cone has a 14-dimensional simplex as its base.
@@
 It is neither an entropic nor an almost entropic point, as was shown in [ZY98].
-The study of the cone of almost entropic points essentially reduces to the question of what neighborhood around the special point is free of almost entropic points, or, equivalently, what neighborhood of the face shared between the Ingleton and non-Ingleton regions is filled with almost entropic points.
+The study of the cone of almost-entropic points essentially reduces to the question of what neighborhood around the special point is free of almost entropic points, or, equivalently, what neighborhood of the face shared between the Ingleton and non-Ingleton regions is filled with almost entropic points.
 [^1] If we denote by $\alpha_{1},\dots,\alpha_{15}$ the convex coordinates in the non-Ingleton simplex (the base simplex of non-Ingleton cone), such that $\alpha_{15}=1$ is the defining equation of the special point, while $\alpha_{15}=0$ is the equation of the face common with the Ingleton-region, then the almost -entropic region is described by the inequality $\alpha_{15}\leq\Phi(\alpha_{1},\dots,\alpha_{14})$ where $\Phi$ is some concave function on the $13$-dimensional simplex.
@@
 The main results of this paper contribute to a better understanding of the behavior of $\Phi$ near the boundary of its domain.
-A detailed discussion of the geometry of the cones $\bar{\Gamma}_{4}^{\star}$, $\Gamma_{4}^{\mathrm{Ing}}$, and $\Gamma_{4}$, and of the non-Ingleton cone can be found, e.g., in [Csi16, Csi25] and in [MC16, MP19].
-In general, the complete description of $\bar{\Gamma}_{4}^{\star}$ remains far from being understood.
-Constraint versions of the Ingleton inequality.
-In what follows we discuss the Ingleton inequality for a four-tuple of random variables.
+Thus, the Ingleton inequality plays a crucial role in the geometry of $\bar{\Gamma}^{*}_{4}$.
+On the one hand, this inequality together with all its permutations provides an inner bound for this cone.
+On the other hand, many non-Shannon type inequalities can be written in the form of the Ingleton inequality with a relatively small additional “correction term”, which give an outer bound for this cone, see the discussion in [WC17, section 2].
+Systematic investigations of the geometry of the cones $\Gamma_{4}$, $\bar{\Gamma}_{4}^{\star}$, and $\Gamma_{4}^{\mathrm{Ing}}$ can be found in [Csi25, MC16, MP20].
+In general, the complete description of $\bar{\Gamma}_{4}^{\star}$ remains far from being understood, despite considerable efforts devoted to this problem.
+1.2 Constraint versions of the Ingleton inequality.
+In what follows we consider the Ingleton inequality in the entropic setting and discuss several versions of this inequality obtained by imposing constraints and/or adding a negligible correction term, that are valid for entropic and almost-entropic points.
 To avoid ambiguity and improve the readability of the notation we denote the four-tuple of jointly distributed random variables $(X,Y,A,B)$, and rewrite (1) with the usual notation for the conditional mutual information as
 $I(X:Y)\leq I(X:Y|A)+I(X:Y|B)+I(A:B)$ (ING)
-We will consider several cases in which some of the quantities
-I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X),I(X:Y|A),I(X:Y|B)I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X),I(X:Y|A),I(X:Y|B) (2)
-are small compared with other information quantities.
-It is known that (ING) holds for joint distributions $(X,Y,A,B)$ satisfying certain independence constraints, that is, whenever some of the quantities in (2) vanish.
-More specifically, the following implications are known:
+When discussing constraint inequalities, we assume that some of the quantities
+I(X:A|Y),I(Y:A|X),I(X:B|Y),I(Y:B|X),I(X:Y|A),I(X:Y|B)I(X:A|Y),I(Y:A|X),I(X:B|Y),I(Y:B|X),I(X:Y|A),I(X:Y|B) (2)
+are small compared to the other information quantities.
+It is known that (ING) holds for joint distributions $(X,Y,A,B)$ satisfying certain independence constraints, that is, whenever some of the quantities in (2) vanish, see [KR13, Stu21].
+We will focus on the following two implications:
 […]
 […]
-as well as the symmetric statements obtained by exchanging $A$ with $B$.
-These conditional constraints were stated explicitly in [KR13], although they were implicitly known earlier, since they follow directly from the infinite families of inequalities established by Matúš in his seminal paper [Mat07].
-We focus on the possible violation of the Ingleton inequality in the regime where some of the mutual information terms in (2) are small but still positive. We consider three cases:
+To the best of our knowledge, the conditional inequalities (3) and (4) were first explicitly formulated in this form in [KR13].
+The proof given in [KR13] relied on a deep and nontrivial result of Matúš, namely the infinite series of conditional inequalities established in [Mat07b].
+Studený proposed in [Stu21] a different proof of (3) and (4), which is more intuitive and technically simpler.
+His argument is based on the following simple lemma.
+Lemma 1 (see lemma 8 in [Stu21]).
+Every discrete distribution of random variables $(X,Y,A)$ satisfying
+I(X:A|Y)=0,I(Y:A|X)=0I(X:A|Y)=0,\ I(Y:A|X)=0
+can be extended by a discrete random variable $W$ such that in the joint distribution $(X,Y,A,W)$ the variable $W$ is functionally dependent on both $X$ and $Y$, and moreover
+$I(X,Y:A|W)=0.$
+The proof of (3) and (4) then proceeds by applying this lemma in combination with standard Shannon inequalities.
+Conceptually, this argument is reminiscent of the classical proof of the Ingleton inequality for linearly representable matroids.
+Remark 2.
+Lemma 1 and its use in proving information inequalities are analogous to the Ahlswede–Körner lemma and its role in the derivation of non-Shannon-type information inequalities in [MMRV02].
+Recall that the Ahlswede–Körner lemma was used in [FKMP18] in computer-assisted proofs of lower bounds for problems of secret sharing.
+We believe that Lemma 1 (and its “smooth” versions discussed below) can be also used in computer-assisted proofs of conditional information inequalities.
+However, the simplicity of Studený’s approach comes at a price.
+Lemma 1 was proved in [Stu21] for entropic points, but not for almost-entropic ones.
+Consequently, the resulting proof of (3) and (4) applies only to entropic points (whereas the more involved argument in [KR13] shows that these inequalities also hold for almost-entropic points).
+In this paper we introduce a technical tool that combines the transparency of Studený’s approach with the power of Matúš’s method.
+We prove a robust version of Lemma 1 that applies to almost-entropic points, see Corollary F and Corollary G.
+Moreover, we establish a trade-off between the error in the assumptions
+$I(X:A|Y)\approx 0,\qquad I(Y:A|X)\approx 0$ (5)
+and the possible deviation in the conclusions
+$H(W|X)\approx 0,\qquad H(W|Y)\approx 0,\qquad I(XY:A|W)\approx 0.$ (6)
+That is, we assume that the approximate equations in (5) hold with precision $\varepsilon$ and analyze the resulting asymptotics in (6).
+Throughout the paper we formulate results of this type using asymptotic notation, without specifying the multiplicative constants hidden in the $O(\cdot)$ terms.
+We believe that refining the asymptotic behavior, rather than optimizing specific constants, will help reveal fundamental properties of the cone of almost-entropic points, see also Remark 3.
+For a discussion of non-asymptotic and unconditional versions of Theorems A, B, and C, see Remark 4.
+Let us mention that (3) and (4) helped to classify the conditional independence structures over $4$ random variables, [Stu21, Boe24] (cf.
+Objective (II) on p.
+(II)).
+The classification of similar structures for $n>4$ remains an open problem, partly due to the rapidly increasing technical difficulties as the dimension grows.
+We hope that the conceptual tools proposed in this paper, such as the robust version of Lemma 1, may contribute to progress in this direction.
+1.3 Main technical results
+We focus on the possible violation of the Ingleton inequality in the regime where some of the mutual information terms in (2) are relatively small but still positive. We consider three cases:
   1. (i)
-    $\max\{I(X:A|Y),\,I(A:Y|X)\}\leq\varepsilon\,I(X:Y)$;
+    $\max\{I(X:A|Y),\,I(Y:A|X)\}\leq\varepsilon\,I(X:Y)$;
   2. (ii)
-    $\max\{I(X:A|Y),\,I(X:Y|A)\}\leq\varepsilon\,I(A:Y)$
+    $\max\{I(X:A|Y),\,I(X:Y|A)\}\leq\varepsilon\,I(Y:A)$
   3. (iii)
-    $\max\{I(A:X|Y),\,I(A:Y|X),\,I(B:X|Y),\,I(B:Y|X)\}\leq\varepsilon\,I(X:Y)$.
+    $\max\{I(X:A|Y),\,I(Y:A|X),\,I(X:B|Y),\,I(Y:B|X)\}\leq\varepsilon\,I(X:Y)$.
 If only two conditional mutual informations are bounded, as in cases (i) and (ii), then the Ingleton inequality holds for such a four-tuple up to an additive error of order $\sqrt{\varepsilon}$.
@@
 Let $(X,Y,A,B)$ and $\varepsilon\geq 0$ be such that
-$\max\left\{I(A:X|Y),I(A:Y|X)\right\}\leq\varepsilon\cdot I(X:Y).$
+$\max\left\{I(X:A|Y),I(Y:A|X)\right\}\leq\varepsilon\cdot I(X:Y).$
 Then
@@
 Let $(X,Y,A,B)$ and $\varepsilon\geq 0$ be such that
-$\max\left\{I(X:Y|A),I(A:Y|X)\right\}\leq\varepsilon\cdot I(X:A).$
+$\max\left\{I(Y:A|X),I(X:Y|A)\right\}\leq\varepsilon\cdot I(X:A).$
 Then
@@
 Let $(X,Y,A,B)$ be a four-tuple of random variables such that
-$\max\{I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X)\}\leq\varepsilon\cdot I(X:Y).$
+$\max\{I(X:A|Y),I(Y:A|X),I(X:B|Y),I(Y:B|X)\}\leq\varepsilon\cdot I(X:Y).$
 Then
 $I(X:Y)\leq I(X:Y|A)+I(X:Y|B)+I(A:B)+O(\varepsilon\cdot\log\varepsilon^{-1})\cdot I(X:Y).$
-Proofs of Theorems A and B are given in Section 4, Theorem C is proven in Section 6.
+These three theorems capture certain geometric features of the boundary of the cone $\bar{\Gamma}^{\star}_{4}$, in particular on how smoothly the curved part of the boundary approaches the flat face defined by (2).
 Remark 3.
@@
 $I(X:Y)$ $\leq I(X:Y|A)+I(X:Y|B)+I(A:B)$
-$\quad{}+\frac{1}{k}I(X:Y|A)+\frac{k+1}{2}\big(I(A:X|Y)+I(A:Y|X)\big)\quad\text{for}\ k\in\mathbb{N}$
+$\quad{}+\frac{1}{k}I(X:Y|A)+\frac{k+1}{2}\big(I(X:A|Y)+I(Y:A|X)\big)\quad\text{for}\ k\in\mathbb{N}$
 and
 $I(X:Y)$ $\leq I(X:Y|A)+I(X:Y|B)+I(A:B)$
-$\quad+\frac{1}{k}I(A:X|Y)+\frac{k+1}{2}\big(I(X:Y|A)+I(A:Y|X)\big)\quad\text{for }k\in\mathbb{N}$
-proven by Matúš in [Mat07] (it suffices to choose a suitable value of $k$).
-In fact, our proof of the lemma above runs parallel to the argument of Matúš — we apply the Copy Lemma and the Ahlswede–Körner lemma, while Matúš used in [Mat07] a similar (and formally equivalent) method of adhesive extensions.
+$\quad+\frac{1}{k}I(Y:A|X)+\frac{k+1}{2}\big(I(X:Y|A)+I(X:A|Y)\big)\quad\text{for }k\in\mathbb{N}$
+proven by Matúš in [Mat07b] (it suffices to choose a suitable value of $k$).
+In fact, our proof of the lemma above runs parallel to the argument of Matúš — we apply the Copy Lemma and the Ahlswede–Körner lemma, while Matúš used in [Mat07b] a similar (and formally equivalent) method of adhesive extensions.
 Likewise, Theorem C follows from the exponential family of inequalities in [DFZ11, Theorem 10],
@@
 by a suitable choice of the parameter $s$.
-This article presents more direct proofs of Theorems ABC without deducing preliminary infinite families of non-Shannon–type inequalities for entropy.
+In this paper we present more direct proofs of Theorems AB, and C without deducing preliminary infinite families of non-Shannon–type inequalities for entropy.
+Another result proved in this paper reveals more subtle properties of probability distributions.
+It can be presented as the following alternative.
+If in a six-tuple of random variables $(X,Y,A,B,C,D)$ the four quantities
+I(X:A|Y),I(Y:A|X),I(X:B|Y),I(Y:B|X)I(X:A|Y),\ I(Y:A|X),\ I(X:B|Y),\ I(Y:B|X)
+are negligibly small, then one of the two properties must be true: a stronger version of the Ingleton inequality for $(X,Y,A,B)$,
+$I(X:Y)\textrm{ is much less than }I(X:Y|A)+I(X:Y|B)+I(A:B),$
+or a slightly weaker version of the Ingleton inequality for $(X,Y,C,D)$,
+$I(X:Y)\ \textrm{ is not much bigger than }I(X:Y|C)+I(X:Y|D)+I(C:D).$
+Here is the exact statement.
+Theorem D (name=,restate=thsubtler).
+If the joint distribution $(X,Y,A,B,C,D)$ satisfies
+$\max\{I(X:A|Y),I(Y:A|X),I(X:B|Y),I(Y:B|X)\}\leq\varepsilon\cdot I(X:Y)$
+Then
+  1. (a)
+    The following alternative holds: for every threshold $0<t<1$ we have a stronger version of Ingleton’s inequality for $(A,B,X,Y)$
+    $I(X:Y)\leq I(X:Y|A)+I(X:Y|B)+I(A:B)+\big(O(\sqrt{\varepsilon})-t\big)I(X:Y)$
+    or a weaker version of Ingleton’s inequality for $(X,Y,C,D)$
+    $I(X:Y)\leq I(X:Y|C)+I(X:Y|D)+I(C:D)+\big(O(\sqrt{\varepsilon})+t\big)I(X:Y)$
+  2. (b)
+    Besides, we have the following inequality for six random variables
+    $\big(2-O(\sqrt{\varepsilon})\big)\cdot I(X:Y)$ $\leq I(X:Y|A)+I(X:Y|B)+I(A:B)$
+    $\;\;+I(X:Y|C)+I(X:Y|D)+I(C:D)$
+Note that the assumptions in the theorem above do not involve variables $C,D$ in any capacity, so that the conclusion of the theorem remains true for arbitrary $C$, $D$.
+While Theorems AB, and C admit proofs based on explicitly derived infinite families of non-Shannon-type inequalities, for Theorem D (proved in Section 5) we are are not aware of any alternative proof that differs substantially from our approach.
+In the next section we give a brief overview of our technique.
+1.4 Methods and ideas
+This article presents more direct proofs of Theorems AB (Section 4), and Theorem C (Section 6) without deducing preliminary infinite families of non-Shannon–type inequalities for entropy.
 This approach offers, in our view, several noteworthy advantages.
 To begin with, it is conceptually more intuitive, as it is grounded in certain structural properties of probability distributions and in the intuition of materializable common information, which we elaborate on below.
-This perspective enables us to reveal and prove properties that are stronger than a mere conditional form of the Ingleton inequality; see Theorem F below.
+This perspective enables us to reveal and prove properties that are stronger than a mere conditional form of the Ingleton inequality, as in Theorem D.
 In addition, the structural property we identify appears particularly amenable to computer-assisted proof systems.
-We believe that this may enable computer-assisted discoveries of new conditional information inequalities, in much the same spirit as non-Shannon–type inequalities and their applications have been established in [DFZ11] and especially in [Far+18, GR19].
-We now outline the “structural property” of a distribution that is at the core of all our proofs.
-Somewhat informally, it states that if $I(A:X|Y)$ and $I(A:Y|X)$ are sufficiently small, then a part of the mutual information between $X$ and $Y$ can be “materialized” in a sense analogous to the notion of Gács–Körner common information [GK73].
-This means that there exists a random variable $W$ that captures a portion of the mutual information between $X$ and $Y$: on the one hand, this $W$ has negligible conditional entropy given either $X$ or $Y$; on the other hand, $W$ retains a significant amount of information: conditioned on $W$, the mutual information between $(X,Y)$ and $A$ becomes very small.
-The lemma below is our main technical tool.
-Lemma D (name=,restate=lWkbounds).
-Suppose a tuple $(X,Y,A,B,\dots)$ is such that for some $\varepsilon>0$
-$\max\left\{I(A:X|Y),I(A:Y|X)\right\}\leq\varepsilon\cdot I(X:Y)$
-Then there exists a sequence of $W_{i}$, $i\in\mathbb{N}_{0}$ such that
+We believe that this may enable computer-assisted discoveries of new conditional information inequalities, in much the same spirit as non-Shannon–type inequalities and their applications have been established in [DFZ11] and especially in [FKMP18, GR19].
+We now outline the structural property of a distribution that is at the core of all our proofs.
+Somewhat informally, it states that if $I(X:A|Y)$ and $I(Y:A|X)$ are sufficiently small, then a part of the mutual information between $X$ and $Y$ can be “materialized” in a sense analogous to the notion of Gács–Körner common information [GK73].
+This means that there exists a random variable $W$ that captures a portion of the mutual information between $X$ and $Y$.
+On the one hand, this $W$ has negligible conditional entropy given either $X$ or $Y$; on the other hand, $W$ retains a significant amount of information—conditioned on $W$, the mutual information between $(X,Y)$ and $A$ becomes very small.
+This idea is formalized in Lemma E below, which is our main technical tool.
+Before Lemma E is formulated, we need to introduce two notions: symmetry among random variables and tropical probability spaces.
+We begin with the symmetry.
+Suppose that a tuple $(X,Y,A,B,Z)$ is such that the variables $A$ and $B$ take value in the same alphabet and such that the joint distribution $p$ of the tuple satisfies
+$p(x,y,a,b,z)=p(x,y,b,a,z)$
+for all samples $x,y,z$ in the alphabets of $X,Y,Z$, respectively, and all $a,b$ in the (shared) alphabet of $A$ and $B$.
+In that case we say that the tuple is symmetric with respect to transposing $A$ and $B$ or variables $A$ and $B$ are symmetric within the tuple.
+Tropical probability spaces were introduced in [MP18].
+Tropicalization associates to a random variable $X$ (or a tuple of random variables) a sequence $X^{tr}:=(X^{(n)})_{n\in\mathbb{N}_{0}}$, whose elements behave, in an appropriate sense, like the joint distribution composed of $n$ independent copies of $X$.
+The entropy of such a tropical random variable is defined as
+$H(X^{tr}):=\lim_{n\to\infty}\tfrac{1}{n}H\big(X^{(n)}\big)$
+with the obvious generalization for tuples and entropic expressions on tuples.
+This is similar to the tensorization trick, which is widely used in analysis, geometry, probability, and related areas.
+In Section 2.2 we discuss the technique of tropicalization in more detail.
+One of the advantages of tropicalization is that any relation between entropic quantities that holds for classical random variables up to a sublinear error holds exactly in the tropical setting.
+From now on we omit the superscript $tr$ and simply refer to such objects as tropical.
+Properties of tuples are understood element-wise along the tropical sequence.
+For example, a tropical pair $(X,Y)$ is called symmetric if the pair $\big(X^{(n)},Y^{(n)}\big)$ is symmetric for all sufficiently large $n$.
+Lemma E (name=,restate=lWkbounds).
+Suppose a tropical tuple $(X,Y,A,B,\dots)$ is such that for some $\varepsilon>0$
+$\max\left\{I(X:A|Y),I(Y:A|X)\right\}\leq\varepsilon\cdot I(X:Y)$
+Then there exists a sequence of tropical random variables $W_{i}$, $i\in\mathbb{N}_{0}$ such that
   1. (a)
@@
 Moreover, if the variables $A$ and $B$ were symmetric in the original distribution $(X,Y,A,B,\ldots)$, then we may assume that this property is preserved for the extended distribution.
-Corollary E (name=,restate=corWbound).
-Suppose a tuple $(X,Y,A,B,\dots)$ is such that for some $\varepsilon>0$
-$\max\left\{I(A:X|Y),I(A:Y|X)\right\}\leq\varepsilon\cdot I(X:Y)$
-Then there exists a $W$ such that
+Corollary F (name=,restate=corWbound).
+Suppose a tropical tuple $(X,Y,A,B,\dots)$ is such that for some $\varepsilon>0$
+$\max\left\{I(X:A|Y),I(Y:A|X)\right\}\leq\varepsilon\cdot I(X:Y)$
+Then there exists a tropical $W$ such that
   1. (a)
@@
     $\max\left\{H(W|X),H(W|Y)\right\}=O(\sqrt{\varepsilon})\cdot I(X:Y)$
-Here an important clarification is in order.
-Lemma D is not literally valid for tuples of jointly distributed random variables $(X,Y,A,B,\dots)$.
-A version of the lemma formulated directly for random variables would necessarily be considerably more cumbersome:
-“For every $\varepsilon^{\prime}>0$ for all sufficiently large $k$, for the scaling of the original distribution $(X,Y,A,B)$ obtained by taking $k$ independent, identically distributed copies of the variables, there exists a variable $W_{i}$ satisfying conditions (a), (b), and (c) within accuracy $\varepsilon^{\prime}\cdot k$.”
-Although this would technically suffice, it would obscure the structural features on which our analysis relies and complicate future integration of the lemma into automated proof frameworks.
-For this reason, Lemma D is stated not for distributions of random variables themselves, but for tropicalized random variables, which are abstract objects derived from ordinary random variables (see [MP20]).
-The definition and basic properties of tropicalized variables are discussed in Section 2.2.
-The tropicalization procedure provides a convenient “interface” that hides certain technical steps — such as scaling up the distribution and controlling sublinear slack terms — and thereby allows us to give a cleaner and more usable formulation of the technical Lemma D and Corollary E.
-By contrast, our main results (Theorems ABC, and F), which rely on Lemma D and Corollary E, can be understood in the more usual way: they apply directly to conventional tuples of random variables.
-We prove Lemma D and Corollary E in Section 3.
+This implies immediately another corollary:
+Corollary G.
+Lemma 1 holds for almost-entropic points.
+The tropicalization procedure provides a convenient “interface” that hides certain technical steps — such as scaling up the distribution and controlling sublinear slack terms — and thereby allows us to give a cleaner and more usable formulation of the technical Lemma E and Corollary F.
+By contrast, our main results (Theorems ABC, and D), which rely on Lemma E and Corollary F, can be understood in the more usual way: they apply directly to conventional tuples of random variables.
+We prove Lemma E and Corollary F in Section 3.
 The proofs may be viewed as compactly capturing the underlying construction originally due to Matúš.
-In particular, the proof of Corollary E requires an a priori unbounded number of applications of the Copy Lemma.
+In particular, the proof of Corollary F requires an a priori unbounded number of applications of the Copy Lemma.
 This fact effectively encodes an infinite family of non-Shannon–type information inequalities, even though they never appear explicitly in the argument.
 At a high level, the proofs of Theorems A and B follow a similar pattern.
-We first apply Corollary E, and then use standard Shannon–type inequalities on the original variables together with the auxiliary variable $W$; see the full argument in Section 4.
+We first apply Corollary F, and then use standard Shannon–type inequalities on the original variables together with the auxiliary variable $W$; see the full argument in Section 4.
 The proofs presented in this paper are entirely conventional in the sense that we explicitly write down the necessary entropy inequalities.
 However, the same arguments can be generated or mechanically verified with computer assistance.
-In practice, one may apply Corollary E in a suitable form, take the properties (a)–(c) for the resulting auxiliary variable $W$, include all basic Shannon-type inequalities involving the variables under consideration (including $W$), and then use a linear-programming solver to deduce the desired conclusion.
-Since the bounds in Corollary E hold only up to $O(\sqrt{\varepsilon})$, the resulting inequality must likewise be understood up to an error term of order $O(\sqrt{\varepsilon})$.
-Furthermore, we combine Lemma D with symmetry considerations and obtain two additional results.
-One of these results demonstrates somewhat subtler properties of tuples of random variables.
-If the six-tuple $(X,Y,A,B,A^{\prime},B^{\prime})$ of random variables satisfies the bounds
-$\max\{I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X)\}\leq\varepsilon\cdot I(X:Y),$
-then the following alternative holds: either a stronger version of the Ingleton inequality for $(X,Y,A,B)$ holds
-$I(X:Y)\textrm{ is much less than }I(X:Y|A)+I(X:Y|B)+I(A:B)$
-or a slightly weaker version of the Ingleton inequality for $(X,Y,A^{\prime},B^{\prime})$ holds
-$I(X:Y)\ \textrm{ is not much bigger than }I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime}).$
-Here is the exact statement.
-Theorem F (name=,restate=thsubtler).
-If the joint distribution $(X,Y,A,B,A^{\prime},B^{\prime})$ satisfies
-$\max\{I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X)\}\leq\varepsilon\cdot I(X:Y)$
-Then
-  1. (a)
-    The following alternative holds: for every threshold $0<t<1$ either we have a stronger version of Ingleton’s inequality for $(A,B,X,Y)$
-    $I(X:Y)\leq I(X:Y|A)+I(X:Y|B)+I(A:B)+\big(O(\sqrt{\varepsilon})-t\big)I(X:Y)$
-    or a weaker version of Ingleton’s inequality for $(A^{\prime},B^{\prime},X,Y)$
-    $I(X:Y)\leq I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})+\big(O(\sqrt{\varepsilon})+t\big)I(X:Y)$
-  2. (b)
-    Besides, we have the following inequality for six random variables
-    $\big(2-O(\sqrt{\varepsilon})\big)\cdot I(X:Y)$ $\leq I(X:Y|A)+I(X:Y|B)+I(A:B)$
-    $\;\;+I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})$
-Note that the assumptions in the theorem above do not involve variables $A^{\prime},B^{\prime}$ in any capacity, so that the conclusion of the theorem remains true for arbitrary $A^{\prime}$, $B^{\prime}$.
-Part (b) of Theorem F probably cannot be proven with more direct approaches (without applying explicitly or implicitly infinitely many instances of the Copy Lemma).
-This theorem does not seem to follow from any previously known results.
-We prove it in Section 5.
-Another result in which we combine Lemma D with symmetry considerations is Theorem C.
-In this theorem we assume that four conditional mutual information terms are of order $\varepsilon$.
-Under this assumption, we obtain a version of the Ingleton bound with an error term of order $O(\varepsilon\log\varepsilon^{-1})$ (instead of $O(\sqrt{\varepsilon})$ in the previous theorems).
-The proof of Theorem C is presented in Section 6.
+In practice, one may apply Corollary F in a suitable form, take the properties (a)–(c) for the resulting auxiliary variable $W$, include all basic Shannon-type inequalities involving the variables under consideration (including $W$), and then use a linear-programming solver to deduce the desired conclusion.
+Since the bounds in Corollary F hold only up to $O(\sqrt{\varepsilon})$, the resulting inequality must likewise be understood up to an error term of order $O(\sqrt{\varepsilon})$.
+In the proof of Theorem C (see Section 6) and Theorem D (see Section 5) we combine Lemma E with symmetry considerations.
 2 Preliminaries
@@
 We denote $n]:=\left\{0,1,\dots,n-1\right\}$ and $2^{[n]}$ will stand for the power set of $\left\{0,1,\dots,n-1\right\}$.
+In this paper, when we talk about random variables, we always mean finitely supported random variable.
 For a collection of jointly distributed random variables $(X_{0},\dots,X_{n-1})$ and $J\in 2^{[n]}$, we denote by $X_{J}$ the joint random variable corresponding to the subcollection $(X_{i}\;{\bm{:}}\;i\in J)$.
-The entropy of a random variable $X$ (respectively, the conditional entropy of $X$ given $Y$) will be denoted $H(X)$ (and, respectively, $H(X|Y)$).
-We use the standard notation for the mutual information in pairs and triples of random variables,
+The entropy of a random variable $X$ will be denoted $H(X)$.
+We use the standard notation for the conditioned entropy, the mutual information in pairs and triples of random variables,
+$H(X|Y):=H(XY)-H(Y)$
 $I(X:Y):=H(X)-H(X|Y),$
@@
 that do not deviate from tensor-powers-sequences too much, more specifically
-$\operatorname{\mathsf{dist}}(\mathbf{X}^{(k)}\otimes\mathbf{X}^{(l)},\mathbf{X}^{(k+l)})\leq C\cdot\varphi(k+l)$
-where $\operatorname{\mathsf{dist}}$ stands for the entropic distance between $n$-tuples of random variables and $\varphi$ is an admissible (sufficiently slowly growing) function, and $C$ is some positive constant that may depend on $\mathbf{X}$.
-For example, we can take $\varphi(t):=t^{3/4}$; see [MP20] for details.
+$\operatorname{\mathsf{dist}}(\mathbf{X}^{(k)}\otimes\mathbf{X}^{(l)},\mathbf{X}^{(k+l)})\leq C\cdot\psi(k+l)$
+where $\operatorname{\mathsf{dist}}$ stands for the entropic distance between $n$-tuples of random variables and $\psi$ is an admissible [^2] A function $\psi$ is admissible if it is monotone and sufficiently slowly growing, more specifically $\int_{1}^{\infty}(\psi(t)/t^{2})\mathrm{d}t<\infty$ function, and $C$ is some positive constant that may depend on the sequence $(\mathbf{X}^{(k)})_{k\in\mathbb{N}}$.
+For example, we can take $\psi(t):=t^{3/4}$ and fix this choice in this article; see [MP20] for details.
 We say that two such sequences $\mathbf{X}^{(k)}=(X_{i}^{(k)}\;{\bm{:}}\;i=0,\dots,n-1)$ and $\mathbf{Y}^{(k)}=(Y_{i}^{(k)}\;{\bm{:}}\;i=0,\dots,n-1)$ are equivalent if
@@
 Entropy of such a tropical random variable is defined as the limit of scaled entropies of individual members of the sequence.
+Other properties of tropical tuples, such as symmetry, are to be understood in the sense that they apply to the tropical sequence element-wise for sufficiently large $n$
+Remark 5.
 Two observations are in order.
-First of all, any linear inequality that is valid for entropies for all tropical random variables implies a similar inequality for entropies of conventional random variables.
-Secondly, if every component of a tropical tuple obeys a given linear inequality with a sublinear error term, e.g., up to $O(1)$, then the tropical tuple itself satisfies this inequality exactly (with no error).
+  1. (i)
+    First of all, any linear inequality that is valid for entropies for all tropical random variables implies the same inequality for entropies of conventional random variables.
+  2. (ii)
+    Secondly, if every component of a tropical tuple obeys a given linear inequality with a sublinear error term, e.g., up to a constant independent of $k$, then the tropical tuple itself satisfies this inequality exactly (with no error).
 Thus, these “tropical” objects can be interpreted as a helpful tool to study properties of the more conventional objects such as probability distributions.
@@
 Thus, we symmetrize the distribution, while changing any $(A,B)$-symmetric entropic expressions by $O(1)$, uniformly with respect to the data.
-Given a tropical tuple we can perform the above procedure elementwise in the sequence of tuples, thereby obtaining exact equalities for the symmetric entropic expressions.
+Given a tropical tuple we can perform the above procedure elementwise in the sequence of tuples, thereby obtaining exact equalities for the symmetric entropic expressions, see Remark 5.
 ∎
-Lemma 3 implies immediately the following corollary.
-Corollary 4.
-Let $(X_{0},\dots,X_{n-1},A,B)$ be a tropical tuple such that for any $J\in 2^{[n]}$ one has $H(X_{J},A)=H(X_{J},B).$ Then there exists another tropical tuple $(X_{0},\dots,X_{n-1},\bar{A},\bar{B})$ such that
-  • the probability distribution of the new tuple is symmetric with respect to the transposition of $\bar{A}$ and $\bar{B}$, and
-  • for any $J\in 2^{[n]}$ one has
-    $H(X_{J},\bar{A})=H(X_{J},A),$
-    $H(X_{J},\bar{B})=H(X_{J},B),$
-    $H(X_{J},\bar{A},\bar{B})=H(X_{J},A,B).$
-Thus, if the entropy profile of a tuple is symmetric, we can replace it with a tuple with the same entropy profile and a symmetric distribution.
-Remark 5.
-Let ${\cal L}_{i}(X,Y,A,B)$, $i=1,2$, be linear combinations of entropies involving $X,Y,A,B$.
-Assume also that these ${\cal L}_{i}$ are symmetric (as formal expressions) with respect to the transposition of the symbols $A$ and $B$, e.g.
-the expression from the Ingleton inequality:
-$\begin{array}[]{l}I(A:B|X)+I(A:B|Y)+I(X:Y)-I(A:B)\\ =H(A,X)+H(B,X)-H(A,B,X)+H(A,Y)+H(B,Y)-H(A,B,Y)\\ \phantom{=}-H(X,Y)-H(A)-H(B)+H(A,B)\end{array}$
-In this setting, if we can prove a statement of the form
-${\cal L}_{1}(X,Y,A,B)=0\Rightarrow{\cal L}_{2}(X,Y,A,B)\geq 0$
-for the distributions symmetric with respect to the transposition of $A$ and $B$, then this implication is true for all distributions.
-Indeed, given an arbitrary distribution $(X,Y,A,B)$, we can apply Lemma 3 and construct its symmetrized version.
-If ${\cal L}_{1}(X,Y,A,B)=0$ for the original distribution, the same equality holds for the symmetrized one (here we use symmetry of ${\cal L}_{1}$).
-So we can conclude that ${\cal L}_{2}(X,Y,A,B)\geq 0$ for the symmetrized distribution.
-Finally, we use symmetry of ${\cal L}_{2}$ and conclude that the same inequality holds for the original distribution.
-This observation allows us in some cases to assume without loss of generality that the random variables under consideration have the property of symmetry.
 2.4 Copy lemma
@@
 can be extended to a longer tuple
-$(X_{i},Y_{j},A^{\prime}\;{\bm{:}}\;i=0,\ldots,k-1;\;j=0,\dots,\ell-1)$
+$(X_{i},Y_{j},A,A^{\prime}\;{\bm{:}}\;i=0,\ldots,k-1;\;j=0,\dots,\ell-1)$
 such that $(X_{0},\ldots,X_{k-1},A)$ and $(X_{0},\ldots,X_{k-1},A^{\prime})$ have the same distribution, and $A^{\prime}$ is independent of $(Y_{0},\ldots,Y_{\ell-1},A)$ conditioned on $(X_{0},\ldots,X_{k-1})$.
@@
 The moreover claim of the lemma follows directly from the definition of the new distribution. ∎
-The entropy profile of the extended tuple $(X_{[k]},Y_{[\ell]},A,A^{\prime})$ is not totally defined by the entropies of the original distribution $(X_{[k]},Y_{[\ell]},A)$.
+The entropy profile of the extended tuple $(X_{[k]},Y_{[\ell]},A,A^{\prime})$ is not totally determined by the entropies of the original distribution $(X_{[k]},Y_{[\ell]},A)$.
 However, some components of the new entropy profile are determined by the definition of $A^{\prime}$: for all $I\subset[k]$ and $J\subset[\ell]$
@@
 2.5 Ahlswede–Körner reduction
-The Ahlswede–Körner reduction, [AK06, AK77], modifies a given tropical $n$-tuple so that some entropies in the new tuple vanish, while some (combinations of) other entropies stay the same.
+The Ahlswede–Körner reduction modifies a given tropical $n$-tuple so that some entropies in the new tuple vanish, while some (combinations of) other entropies stay the same.
 Lemma 8 (Ahlswede–Körner reduction).
-Let us consider a tropical $n$-tuple $(X_{i}:i=0,\dots,n-1)$ and a subcollection $(X_{i}:i=0,\dots,k-1)$ of it, where $k\leq n$.
+Let us consider a tropical $n$-tuple $(X_{i}:i=0,\dots,n-1)$ and a subcollection $(X_{i}:i=0,\dots,k-1)$ of it, where $k<n$.
 Then there exists an $(n+1)$-tuple $(X_{0},\dots,X_{n-1},X^{\prime}_{n-1})$, such that
@@
 $X^{\prime}_{n-1}:=\operatorname{\mathsf{AK}}(X_{n-1}\mid X_{0},\dots,X_{k-1}).$
+Remark 9.
+The claim $H(X_{J}|X^{\prime}_{n-1})=H(X_{J}|X_{n-1})$ in the Ahlswede–Körner reduction can be reformulated as $I(X_{J}:X^{\prime}_{n-1})=I(X_{J}:X_{n-1})$.
+Remark 10.
+The extension of the distribution guaranteed by the Ahlswede–Körner reduction is not unique, i.e., there may exist different (nonisomorphic) extensions of $(X_{i}:i=0,\dots,n-1)$ satisfying the conditions of Lemma 8.
+We abuse notation and write $X^{\prime}_{n-1}:=\operatorname{\mathsf{AK}}(X_{n-1}\mid X_{0},\dots,X_{k-1})$, meaning that any such extension may be chosen.
+Essentially Lemma 8 claims that the variable $X_{n-1}$ can be “reduced” to a new random variable $X^{\prime}_{n-1}$, that will be functionally dependent on $X_{[k]}$, while all the quantities $I(X_{J}:X_{n-1})$, $J\in 2^{[k]}$, will remain preserved after the reduction.
+In Fig.
+1 we illustrate this reduction for the case $k=n=2$.
+This statement is not literally true for untropicalized tuple, but only hold up to an additive error, which is sublinear with respect to $H(X_{[n]})$.
+Recall that tropical tuple is a growing sequence of the classical tuples of random variables such that entropic quantities of individual members of the sequence grow quasi-linearly.
+Entropic quantity evaluated on the tropical tuple is the limit of the corresponding quantities of individual tuples normalized by the index in the sequence.
+Thus after tropicalization sublinear errors vanish.
+[…]
+Figure 1: Diagrams representing the Ahlswede–Körner reduction for $k=2$.
+The reduced variable $X_{2}^{\prime}=\operatorname{\mathsf{AK}}(X_{n-1}\mid X_{0},X_{1})$ has the same mutual information with $X_{0}$, $X_{1}$, and $(X_{0},X_{1})$ as the original variable $X_{2}$; at the same time, $H(X_{2}^{\prime}|X_{0},X_{1})$ vanishes.
+History of the Ahlswede–Körner reduction.
+This technical tool goes back to the work of Ahlswede and Körner in the 1970s [AK06, AK77].
+For readers interested in a proof of this lemma, we can point to several sources.
+For the case $k=2$ and $n=3$, the statement is proven in [MMRV02, Lemma 5], and the argument extends straightforwardly to arbitrary $k$ and $n$.
+A similar argument can be found in [MP19, Theorem 3.1].
+The latter article, however, uses somewhat different terminology.
+The lemma can also be obtained as a special case of more general techniques.
+Namely, one may first introduce a copy $X^{\prime}_{n-1}$ of $X_{n-1}$ (preserving its joint distribution with $(X_{0},\dots,X_{k-1})$), and then perform tightening as in [MC16, Lemma 3] (with $L=\{X^{\prime}_{n-1}\}$ and $t=H(X_{n-1}\mid X_{0},\dots,X_{k-1})$).
+The technically most involved part of this argument is established in [Mat07a, Theorem 3], which can be viewed as a generalization of the Slepian–Wolf theorem [SW73].
+An overview of the tightening technique and its applications can be found in [Csi25].
 3 Main construction
-In this section we prove our main technical results, Lemma D and Corollary E, which are repeated below verbatim.
+In this section we prove our main technical results, Lemma E and Corollary F, which are repeated below verbatim.
 They are systematically applied in all proofs of the article.
@@
 Step 1 (duplication):
-First of all, we apply the Copy Lemma $k$ times to $A$ under $(X,Y)$ and obtain new variables $A^{\prime}_{1},\ldots,A^{\prime}_{k}$ such that
-  • for each $i=1,\ldots,k$ the conditional distribution of $A^{\prime}_{i}$ given $(X,Y)$ is the same as the conditional distribution of $A$;
+First of all, we apply the Copy Lemma $k$ times to $A$ under $(X,Y)$ and obtain new variables $A^{\prime}_{1},\ldots,A^{\prime}_{k}$ such that for each $i=1,\ldots,k$
+  • the new variable $A^{\prime}_{i}$ has the same distribution as $A$ conditional on $(X,Y)$;
+  • conditional on $(X,Y)$, the new variable $A^{\prime}_{i}$ is independent of $A^{\prime}_{1},\ldots,A^{\prime}_{i-1},A,B,C_{0},\dots,C_{n-1}$.
+Observe that
+$I(A^{\prime}_{1},\ldots,A^{\prime}_{k}$ $:A,B,C_{0},\dots,C_{n-1}|X,Y)$
+$=\sum\limits_{i=1}^{k}I(A^{\prime}_{i}:A,B,C_{0},\dots,C_{n-1}|X,Y,A^{\prime}_{1},\ldots,A^{\prime}_{i-1})$
+$\leq\sum\limits_{i=1}^{k}I(A^{\prime}_{i}:A^{\prime}_{1},\ldots,A^{\prime}_{i-1},A,B,C_{0},\dots,C_{n-1}|X,Y)=0$
+(the first step is the chain rule for the mutual information; the second step is an instance of the Shannon-type inequality $I(U:V|W)\leq I(U:V,W)$; the last equality follows from the definition of $A^{\prime}_{i}$ in the Copy Lemma).
+Thus, we conclude that
   • the random variables $A^{\prime}_{1},\ldots,A^{\prime}_{k}$ and the tuple $(A,B,C_{0},\ldots,C_{n-1})$ are all together mutually independent conditional on $(X,Y)$.
@@
 Step 2 (reduction):
-Further, we apply the Ahlswede–Körner reduction to $\tilde{A}_{k}$ under $(A,B,C_{0},\ldots,C_{n-1})$ and denote the result by
+Further, we apply the Ahlswede–Körner reduction to $\tilde{A}_{k}$ under the tuple $(A,B,C_{0},\ldots,C_{n-1})$ and denote the result by
 $W_{k}:=\operatorname{\mathsf{AK}}(\tilde{A}_{k}\mid X,Y,A,B,C_{0},\ldots,C_{n-1}).$
@@
 $H(A,W_{k})=H(B,W_{k}),\quad H(A,C_{J},W_{k})=H(B,C_{J},W_{k}),\text{ etc}.$
-3.2 Proof of Lemma D
-We use $W_{[k]}$ constructed in the previous section to prove the lemma.
-To prove assertion (a), we note that by the Copy Lemma, the variable $\tilde{A}_{k}$ is independent of $(A,B,C_{0},\ldots,C_{n-1})$ conditioned on $(X,Y)$.
-The same holds for $W_{k}=\operatorname{\mathsf{AK}}(\tilde{A}_{k}|X,Y,A,B,C_{0},\dots,C_{n-1})$, by the definition of the Ahlswede–Körner reduction.
-Thus we have
-$H(W_{k}|X,Y)=H(W_{k}|X,Y,A,B,C_{0},\dots,C_{n-1})=0$
-To prove assertion (b) it suffices to observe that the conditional distributions of $(X,Y)$ given $(A^{\prime}_{1},\ldots,A^{\prime}_{k+1})$ and given $(A^{\prime}_{1},\ldots,A^{\prime}_{k},A)$ are identical.
-To prove (c), we observe that
-$H(W_{k}|X)$ $=I(\tilde{A}_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1}|X)$
+3.2 Proof of Lemma E
+We recall that we define each variable $A^{\prime}_{i+1}$ with help of the Copy Lemma so that
+  1. (a)
+    it has the same distribution as $A$ conditional on $(X,Y)$;
+  2. (b)
+    $A^{\prime}_{i+1}$ and $(A^{\prime}_{1},\ldots,A^{\prime}_{i},A,B,C_{0},\ldots,C_{n-1})$ are independent conditional on $(X,Y)$.
+Claim 11.
+For each $k$
+$(A,B,C_{0},\ldots,C_{n-1})\to(X,Y)\to(A^{\prime}_{1},\ldots,A^{\prime}_{k})$
+is a Markov chain.
+This lemma follows easily from the definition of the quantities $A^{\prime}_{i}$. For completeness of exposition, we give a detailed derivation in terms of information inequalities.
+Proof of the claim:.
+The argument is based on the property (b) above with $i=1,\ldots,k-1$. We use the chain rule, and then apply $i$ times inequalities having the form (23) and obtain
+$I(A,B,C_{0},\ldots,C_{n-1}:A^{\prime}_{1},\ldots,A^{\prime}_{i}|X,Y)$ $=\sum\limits_{j=0}^{i}I(A,B,C_{0},\ldots,C_{n-1}:A^{\prime}_{j}|X,Y,A^{\prime}_{1},\ldots,A^{\prime}_{j-1})$
+$\leq\sum\limits_{j=0}^{i}I(A^{\prime}_{1},\ldots,A^{\prime}_{j-1},A,B,C_{0},\ldots,C_{n-1}:A^{\prime}_{j}|X,Y)$
+$=0,$
+and the claim is proven. ∎
+Now we can prove assertion (a) of Lemma E. Claim 11 implies
+$I(\tilde{A}_{k}:A,B,C_{0},\ldots,C_{n-1},X,Y)$ $\leq I(\tilde{A}_{k}:X,Y)+I(\tilde{A}_{k}:A,B,C_{0},\ldots,C_{n-1}|X,Y)$
+$=I(\tilde{A}_{k}:X,Y).$
+From the definition of the Ahlswede–Körner reduction (see Remark 9) it follows
+$I(W_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1})$ $=I(\tilde{A}_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1})$
+$\leq I(\tilde{A}_{k}:X,Y)=I(W_{k}:X,Y).$
+Therefore, with the chain rule we obtain
+$I(W_{k}:A,B,C_{0},\ldots,C_{n-1}|X,Y)$ $=I(W_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1})-I(W_{k}:X,Y)$
+$=0.$
+Hence,
+$H(W_{k}|X,Y)$ $=H(W_{k}|X,Y,A,B,C_{0},\ldots,C_{n-1})+I(W_{k}:A,B,C_{0},\ldots,C_{n-1}|X,Y)$
+$=H(W_{k}|X,Y,A,B,C_{0},\ldots,C_{n-1})=0$
+(the last equality follows from the Ahlswede–Körner reduction).
+To prove assertion (b) of Lemma E, we observe that the following conditional distributions are all identical:
+  1. (i)
+    distribution of $A$ given the value of $(X,Y,A^{\prime}_{1},\ldots,A^{\prime}_{i})$
+  2. (ii)
+    distribution of $A$ given the value of $(X,Y)$
+    (this distribution is equivalent to the previous one due to the Markov property)
+  3. (iii)
+    distribution of $A^{\prime}_{i+1}$ given the value of $(X,Y)$
+    (equivalent to the previous one due to the property (a) above)
+  4. (iv)
+    distribution of $A^{\prime}_{i+1}$ given the value of $(X,Y,A^{\prime}_{1},\ldots,A^{\prime}_{i})$
+    (equivalent to the previous one due to the property (b) above).
+Comparing (i) and (iv), we conclude that the conditional distributions
+[…]
+and
+[…]
+are the same. It follows that
+$\left\{\begin{array}[]{l}H(X,Y|\tilde{A}_{i+1})=H(X,Y|\tilde{A}_{i},A)=H(X,Y,A|\tilde{A}_{i})-H(A|\tilde{A}_{i}),\\ H(X|\tilde{A}_{i+1})=H(X|\tilde{A}_{i},A)=H(X,A|\tilde{A}_{i})-H(A|\tilde{A}_{i}),\\ H(Y|\tilde{A}_{i+1})=H(Y|\tilde{A}_{i},A)=H(Y,A|\tilde{A}_{i})-H(A|\tilde{A}_{i}).\end{array}\right$.
+The Ahlswede–Körner reduction guarantees that
+[…]
+and
+$\left\{\begin{array}[]{l}H(X,Y|\tilde{A}_{i})=H(X,Y|W_{i}),\\ H(X|\tilde{A}_{i})=H(X|W_{i}),\\ H(Y|\tilde{A}_{i})=H(Y|W_{i}),\\ H(X,Y,A|\tilde{A}_{i})=H(X,Y,A|W_{i}),\\ H(X,A|\tilde{A}_{i})=H(X,A|W_{i}),\\ H(Y,A|\tilde{A}_{i})=H(Y,A|W_{i}),\\ H(A|\tilde{A}_{i})=H(A|W_{i}).\end{array}\right$.
+We combine these observations and obtain
+$I(X:Y|W_{i+1})$ $=H(X|W_{i+1})+H(Y|W_{i+1})-H(X,Y|W_{i+1})$
+$=H(X|\tilde{A}_{i+1})+H(Y|\tilde{A}_{i+1})-H(X,Y|\tilde{A}_{i+1})$
+$=H(X|\tilde{A}_{i},A)+H(Y|\tilde{A}_{i},A)-H(X,Y|\tilde{A}_{i},A)$
+$=H(X,A|\tilde{A}_{i})+H(Y,A|\tilde{A}_{i})-H(X,Y,A|\tilde{A}_{i})-H(A|\tilde{A}_{i})$
+$=H(X,A|W_{i})+H(Y,A|W_{i})-H(X,Y,A|W_{i})-H(A|W_{i})$
+$=I(X:Y|W_{i},A),$
+and (b) is proven.
+To prove (c) of Lemma E, we observe that
+$H(W_{k}|X)$ $=H(W_{k},X)-H(X)$
+$=H(X|W_{k})+H(W_{k})-H(X)$
+$=H(X|\tilde{A}_{k})+I(\tilde{A}_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1})-H(X)$
 /* from the AK Lemma */
-$=I(\tilde{A}_{k}:X,Y|X)$
-/* since $\tilde{A}_{k}\to(XY)\to(A,B,C_{0},\ldots,C_{n-1})$ is a Markov chain */
+$=I(\tilde{A}_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1})-I(\tilde{A}_{k}:X)$
+$=I(\tilde{A}_{k}:X,Y)-I(\tilde{A}_{k}:X)$
+/* from the Copy Lemma: $\tilde{A}_{k}$ is independent of
+$(A,B,C_{0},\ldots,C_{n-1})$ conditioned on $(X,Y)$ */
 $=I(\tilde{A}_{k}:Y|X)=H(\tilde{A}_{k}|X)-H(\tilde{A}_{k}|X,Y).$
@@
 Thus, $H(W_{k}|X)=H(\tilde{A}_{k}|X)-H(\tilde{A}_{k}|X,Y)$ rewrites to
-$H(W_{k}|X)=k\cdot\left(H(A|X)-H(A|X,Y)\right)=k\cdot I(A:Y|X)\leq k\cdot\varepsilon\cdot I(X:Y).$
+$H(W_{k}|X)\leq k\cdot\left(H(A|X)-H(A|X,Y)\right)=k\cdot I(Y:A|X)\leq k\cdot\varepsilon\cdot I(X:Y).$
 A similar argument gives the bound for $H(W_{k}|Y)$.
@@
 We already know that the entropy profile $(X,Y,A,B,C_{0},\ldots,C_{n-1},W_{k})$ is symmetric under swapping $A$ and $B$.
-To make the property stronger and enforce the symmetry of the tropical tuple, we apply Corollary 4.
-3.3 Proof of Corollary E
+Applying Lemma 6 to the distribution proves the claim.
+3.3 Proof of Corollary F
 […]
 Proof.
-The argument is based on Lemma D, where we will choose a suitable $k$ and let $W=W_{k}$. We apply Lemma D with the minimal $k$ such that
+The argument is based on Lemma E, where we will choose a suitable $k$ and let $W=W_{k}$. We apply Lemma E with the minimal $k$ such that
 $I(X:Y|W_{k})\leq I(X:Y|A,W_{k})+\delta\cdot I(X:Y)$
@@
 $I(X:Y|W_{i-1})-I(X:Y|W_{i})$
-is at least $\delta\cdot I(X:Y)$. Since the sum of all such gaps does not exceed $I(X:Y)$, the number of steps $k$ is at most $k\leq 1/\delta.$ From Lemma D(c) we get
+is at least $\delta\cdot I(X:Y)$. Since the sum of the gaps
+$\big(I(X:Y)-I(X:Y|W_{1})\big)$ $+\big(I(X:Y|W_{1})-I(X:Y|W_{2})\big)+\ldots$
+$\ldots+\big(I(X:Y|W_{k-1})-I(X:Y|W_{k})\big)=I(X:Y)-I(X:Y|W_{k})$
+does not exceed $I(X:Y)$, the number of steps $k$ is at most $k\leq 1/\delta.$ From Lemma E(c) we get
 $\max\left\{H(W_{k}|X),H(W_{k}|Y)\right\}\leq k\cdot\varepsilon\cdot I(X:Y)\leq\frac{\varepsilon}{\delta}\cdot I(X:Y)$
 It follows that
-$I($ $A:X,Y|W_{k})=I(A:X|Y,W_{k})+I(A:Y|X,W_{k})+I(A:X:Y|W_{k})$
-$\leq I(A:X|Y)+H(W_{k}|X)+I(A:Y|X)+H(W_{k}|Y)+I(A:X:Y|W_{k})$
-$\leq I(A:X|Y)+I(A:Y|X)+I(A:X:Y|W_{k})+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
-$\leq I(A:X:Y|W_{k})+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
+$I($ $X,Y:A|W_{k})=I(X:A|Y,W_{k})+I(Y:A|X,W_{k})+I(X:Y:A|W_{k})$
+$=I(X:A|Y,W_{k})+I(Y:A|X,W_{k})+I(X:Y:A|W_{k})$
+$\leq I(X:A|Y)+H(W_{k}|Y)+I(Y:A|X)+H(W_{k}|X)+I(X:Y:A|W_{k})$
+(here we used twice (27) with a suitable substitution of variables)
+$\leq I(X:A|Y)+I(Y:A|X)+I(X:Y:A|W_{k})+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
+$\leq I(X:Y:A|W_{k})+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
 $\leq I(X:Y|W_{k})-I(X:Y|A,W_{k})+O\left(\frac{\varepsilon}{\delta}\right)\cdot I(X:Y)$
@@
 In this section we prove that if the conditional information quantities
-I(A:X|Y),I(A:Y|X)I(A:X|Y),\ I(A:Y|X)
+I(X:A|Y),I(Y:A|X)I(X:A|Y),\ I(Y:A|X)
 are of order $\varepsilon$, then a version of the Ingleton inequality (ING) is valid with a small correction term of order $\sqrt{\varepsilon}$; see Theorem A.
@@
 Proof.
-First of all, we apply Corollary E and take a $W$ such that
+First of all, we apply Corollary F and take a $W$ such that
   • $H(W|X)=O(\sqrt{\varepsilon})\cdot I(X:Y)$,
   • $H(W|Y)=O(\sqrt{\varepsilon})\cdot I(X:Y)$,
-  • $I(A:X,Y|W)=O(\sqrt{\varepsilon})\cdot I(X:Y)$.
+  • $I(X,Y:A|W)=O(\sqrt{\varepsilon})\cdot I(X:Y)$.
 The rest of the proof is an application of suitable Shannon–type inequalities and can therefore be verified with the help of any solver that handles Shannon–type entropy inequalities.
@@
 see inequality (28) in the Appendix. For the $W$ chosen above, this inequality implies
-$\begin{split}I(X:Y|A)&\geq H(W|A)+I(X:Y|A,W)-O(\sqrt{\varepsilon})\cdot I(X:Y)\\ &\geq H(W|A)+I(X:Y|W)-O(\sqrt{\varepsilon})\cdot I(X:Y).\end{split}$
+$\begin{split}I(X:Y|A)&\geq H(W|A)+I(X:Y|A,W)-O(\sqrt{\varepsilon})\cdot I(X:Y)\\ &\geq H(W|A)+I(X:Y|W)-I(X:Y:A|W)-O(\sqrt{\varepsilon})\cdot I(X:Y)\\ &\geq H(W|A)+I(X:Y|W)-I(X,Y:A|W)-O(\sqrt{\varepsilon})\cdot I(X:Y)\\ &\geq H(W|A)+I(X:Y|W)-O(\sqrt{\varepsilon})\cdot I(X:Y).\end{split}$
 It follows that
@@
 Proof.
-Similarly to the previous argument, we apply Corollary E (switching $A$ and $Y$ in the assumptions and the conclusion of the lemma) and take a $W$ such that
+Similarly to the previous argument, we apply Corollary F (switching $A$ and $Y$ in the assumptions and the conclusion of the lemma) and take a $W$ such that
   • $H(W|X)=O(\sqrt{\varepsilon})\cdot I(X:A)$,
@@
 First, applying several times the chain rule for the mutual information, we obtain
-$\begin{split}I(X:Y)&\leq I(X:Y)+I(W:Y|X)\\ &=I(X,W:Y)\\ &=I(W:Y)+I(X:Y|W),\\ &\leq I(W:Y)+I(X:Y|W)+I(A:Y|X,W)\\ &=I(W:Y)+I(X,A:Y|W)\\ &=I(W:Y)+O(\Delta),\end{split}$
+$\begin{split}I(X:Y)&\leq I(X:Y)+I(Y:W|X)\\ &=I(X,W:Y)\\ &=I(Y:W)+I(X:Y|W),\\ &\leq I(Y:W)+I(X:Y|W)+I(Y:A|X,W)\\ &=I(Y:W)+I(X,A:Y|W)\\ &=I(Y:W)+O(\Delta),\end{split}$
 where the last step uses the assumption $I(X,A:Y|W)=O(\Delta)$.
@@
 5 Subtler claims with the $O(\sqrt{\varepsilon})$ precision
-In this section we combine Lemma D with symmetry considerations, derive a new information inequality for six-tuples of random variables, and prove Theorem F from the introduction.
+In this section we combine Lemma E with symmetry considerations, derive a new information inequality for six-tuples of random variables, and prove Theorem D from the introduction.
 […]
+Remark 15.
+The constant hidden in the $O(\cdot)$-notation does not depend on $t$ nor on the distribution of $(X,Y,A,B,C,D)$.
 Proof.
-First of all, we observe that the inequalities in both claims (a) and (b) are symmetric as syntactic expressions with respect to the transposition of $A$ and $B$.
-Therefore, we can apply Lemma 3 (see also Remark 5) and assume without loss of generality that $A$ and $B$ are symmetric in the distribution $(X,Y,A,B,A^{\prime},B^{\prime})$.
-Now we are ready to prove the theorem.
-To address the first claim of the theorem, we apply Corollary E to construct $W$ such that
+We first show that, without loss of generality, we may assume that the distribution of the tuple $(X,Y,A,B,C,D)$ is symmetric with respect to transposition of $A$ and $B$.
+We begin with tropicalization: we replace the original tuple by its tropical version $(X,Y,A,B,C,D)^{tr}$, where each variable is replaced by the joint of $n$ independent copies of itself.
+The assumptions of the theorem still hold for the tropical tuple; see Remark 5.
+We then apply Lemma 6 and transform this tropical tuple into a symmetric one that still satisfies the assumptions of the theorem.
+Note that the entropies of the symmetrized tropical tuple coincide with those of the original tuple.
+Thus, if the theorem holds for the symmetrized tuple, then it also holds for the original one.
+Now we are ready to prove the theorem. To address the first claim of the theorem, we apply Corollary F to construct $W$ such that
   • $H(W|X)=O(\sqrt{\varepsilon})\cdot I(X:Y)$,
   • $H(W|Y)=O(\sqrt{\varepsilon})\cdot I(X:Y)$,
-  • $I(A:X,Y|W)=O(\sqrt{\varepsilon})\cdot I(X:Y)$.
-Due to Lemma D (the moreover part) we may assume that all entropy quantities for $(X,Y,A,W)$ are the same as the corresponding quantities for $(X,Y,B,W)$. Similarly to (5), using the symmetry we have
+  • $I(X,Y:A|W)=O(\sqrt{\varepsilon})\cdot I(X:Y)$.
+Due to Lemma E (the moreover part) we may assume that all entropy quantities for $(X,Y,A,W)$ are the same as the corresponding quantities for $(X,Y,B,W)$. Similarly to (7), using the symmetry we have
 $H(W|A)$ $\leq I(X:Y|A)-I(X:Y|W)+O(\sqrt{\varepsilon})\cdot I(X:Y)$
@@
 This gives the first part of the alternative in claim (a).
-On the other hand, for all $(X,Y,A^{\prime},B^{\prime},W)$ we can follow the usual scheme of the proof of the Ingleton’s inequality with $W$ representing the common information of $X$ and $Y$.
+On the other hand, for all $(X,Y,C,D,W)$ we can follow the usual scheme of the proof of the Ingleton’s inequality with $W$ representing the common information of $X$ and $Y$.
 More formally, we add Shannon–type inequalities
-$H(W)$ $\leq H(W|A^{\prime})+H(W|B^{\prime})+I(A^{\prime}:B^{\prime}),$
-$H(W|A^{\prime})$ $\leq H(W|X)+H(W|Y)+I(X:Y|A^{\prime}),$
-$H(W|B^{\prime})$ $\leq H(W|X)+H(W|Y)+I(X:Y|B^{\prime}),$
+$H(W)$ $\leq H(W|C)+H(W|D)+I(C:D),$
+$H(W|C)$ $\leq H(W|X)+H(W|Y)+I(X:Y|C),$
+$H(W|D)$ $\leq H(W|X)+H(W|Y)+I(X:Y|D),$
 $I(X:Y)$ $\leq H(W)+I(X:Y|W)$
 to obtain
-$I(X:Y)$ $\leq I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})$
+$I(X:Y)$ $\leq I(X:Y|C)+I(X:Y|D)+I(C:D)$
 $+2\big(H(W|X)+H(W|Y)\big)+I(X:Y|W)$
 Therefore, for the chosen $W$ we have
-$\begin{split}I(X:Y)&\leq I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})\\ &+O(\sqrt{\varepsilon})\cdot I(X:Y)+I(X:Y|W)\end{split}$ (10)
+$\begin{split}I(X:Y)&\leq I(X:Y|C)+I(X:Y|D)+I(C:D)\\ &+O(\sqrt{\varepsilon})\cdot I(X:Y)+I(X:Y|W)\end{split}$ (12)
 Thus, if $t\geq I(X:Y|W)/I(X:Y)$, we obtain
-$I(X:Y)\leq I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})+\big(O(\sqrt{\varepsilon})+t\big)\cdot I(X:Y).$
+$I(X:Y)\leq I(X:Y|C)+I(X:Y|D)+I(C:D)+\big(O(\sqrt{\varepsilon})+t\big)\cdot I(X:Y).$
 This gives the second part of the alternative.
@@
 Proof.
+The conclusion of the theorem (the Ingleton inequality) is trivial in case $I(X:Y)=0$.
+So, in what follows we assume without loss of generality that the mutual information $I(X:Y)$ is strictly positive.
 Since the inequalities in both the assumption and the conclusion of the theorem are symmetric with respect to swapping $A$ and $B$, we may apply Lemma 6 to the four-tuple $(X,Y,A,B)$.
 Thus, without loss of generality, we may assume that the distribution is symmetric (see also Remark 5).
-Now we use Lemma D and symmetry of the distribution to find a sequence of extensions $(X,Y,A,B,W_{i})$, $i=0,1,\dots$ satisfying
+Now we use Lemma E and symmetry of the distribution to find a sequence of extensions $(X,Y,A,B,W_{i})$, $i=0,1,\dots$ satisfying
 $H(W_{i}|X,Y)$ $=0$
@@
 Then
-$\delta\cdot I(X:Y)\leq I(X:Y|W_{k})\leq\lambda I(X:Y|W_{k-1})\leq\ldots\leq\lambda^{k}\cdot I(X:Y)$
-Therefore
+$\delta\cdot I(X:Y)\leq I(X:Y|W_{k})\leq\lambda I(X:Y|W_{k-1})\leq\ldots\leq\lambda^{k}\cdot I(X:Y).$
+Since we assume that $I(X:Y)>0$, we may conclude that
 $k\leq\frac{\ln\delta}{\ln\lambda}$ (15)
@@
 6.2 Case 2: $I(X:Y|W_{k})\leq\delta\cdot I(X:Y)$
-We apply inequality (25) from the Appendix, and use the properties of $W_{i}$ from Lemma D to estimate
+We apply inequality (29) from the Appendix, and use the properties of $W_{i}$ from Lemma E to estimate
 $I(X:Y)$ $\leq I(X:Y|A)+I(X:Y|B)+I(A:B)$
-$\quad+I(X:Y|W_{k})+I(X:W_{k}|Y)+I(W_{k}:Y|X)$
+$\quad+I(X:Y|W_{k})+I(X:W_{k}|Y)+I(Y:W_{k}|X)$
 $\leq I(X:Y|A)+I(X:Y|B)+I(A:B)+(\delta+2k\cdot\varepsilon)I(X:Y)$
@@
 $I(X,Y:A|B)=I(X:A|B)+I(Y:A|X,B).$
+Observe that it implies
+$I(X,Y:A|B)\geq I(Y:A|X,B).$ (23)
 Besides, we apply several times the well-known Shannon–type inequality
@@
 $H(Z|W)\leq H(Z|X,W)+H(Z|Y,W)+I(X:Y|W),$ (25)
-In addition, we use three other Shannon–type inequalities:
+In addition, we use several other Shannon–type inequalities. First of all, we use
 $I(X:Y)\leq H(W)+I(X:Y|W)$ (26)
-(which is equivalent to the sum of $H(W|X)\geq 0$ and $I(W:X|Y)\geq 0$),
+(which is equivalent to the sum of $H(W|X)\geq 0$ and $I(X:W|Y)\geq 0$) and its conditional version
+$I(X:Y|Z)\leq H(W|Z)+I(X:Y|W,Z).$ (27)
+Besides, we need the inequality
 $H(W)+I(X:Y|W)\leq I(X:Y)+\big(H(W|X)+H(W|Y)\big)$
@@
 $H(W|A)+I(X:Y|W,A)\leq I(X:Y|A)+\big(H(W|X,A)+H(W|Y,A)\big).$ (28)
-Finally, we use a version of the classical non-Shannon–type inequality from [ZY98] in the form with five variables (proven in [Mak+02]),
+Finally, we use a version of the classical non-Shannon–type inequality from [ZY98] in the form with five variables (proven in [MMRV02, theorem 1 for $n=2$]),
+$H(A,B)+2I(X:Y:W)\leq I(X:Y|A)+I(X:Y|B)+H(A)+H(B)+I(XY:W).$
+This inequality is equivalent to
+$\begin{split}2I(X:Y:W)&\leq I(X:Y|A)+I(X:Y|B)+I(A:B)\\ &\quad+\left(I(X:Y:W)+I(X:W|Y)+I(Y:W|X)\right).\end{split}$
+Since $I(X:Y)=I(X:Y:W)+I(X:Y|W)$, it can be rewritten in the form convenient for our application:
 $\begin{split}I(X:Y)&\leq I(X:Y|A)+I(X:Y|B)+I(A:B)\\ &\quad+I(X:Y|W)+I(X:W|Y)+I(Y:W|X).\end{split}$ (29)